### v1

# Holonomy of the Obata connection on 2-step hypercomplex nilmanifolds

Abstract
We study the holonomy of the Obata connection on 2-step hypercomplex nilmanifolds. By explicitly computing the curvature tensor, we determine the conditions under which the Obata connection is flat, showing that this depends on the nilpotency step of each complex structure. In particular, we show that the holonomy algebra of the Obata connection is always an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$: to establish this, we prove the $\mathbb{H}$-solvable conjecture for 2-step hypercomplex nilmanifolds. Furthermore, we provide new examples of $k$-step nilpotent hypercomplex nilmanifolds, with arbitrary $k$, which are not Obata flat.
1 Introduction
Nilmanifolds, that is, compact quotients $\G\Gamma\backslash$ G of connected and simply connected nilpotent Lie groups $G$ by discrete subgroups $\Gamma$, have played a fundamental role in differential geometry and topology since the seminal works of Mal’cev and Mostow. They provide a natural class of spaces where algebraic and geometric properties interact tightly, offering a fair balance between algebraic manageability and rich geometric behavior. In low dimensions, the latter allows classification results for various geometric structures (e.g., [Sal, FPS, Ug]), especially when these are left-invariant.
Invariant complex structures have been particularly investigated, largely due to the foundational work of Salamon [Sal], where it is shown that the existence of invariant complex structures $J$ on nilmanifolds admits special features: e.g., the existence of a so-called Salamon basis [Sal, Theorem 1.3], and the fact that the subspace $\mathfrak{g},\mathfrak{g}]+J[\mathfrak{g},\mathfrak{g}]$ is always a proper ideal of the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ [Sal, Corollary 1.4].
In the hypercomplex setting, many of these nice behaviors fail to carry over in a unified way. Each complex structure $J_{\alpha}$ ($\alpha=1,2,3$) may satisfy suitable compatibility conditions individually, but it is not clear whether these structures collectively induce analogous global properties. For instance, it is natural to ask whether the subspace
$\mathfrak{g}_{1}^{\mathbb{H}}:=[\mathfrak{g},\mathfrak{g}]+J_{1}[\mathfrak{g},\mathfrak{g}]+J_{2}[\mathfrak{g},\mathfrak{g}]+J_{3}[\mathfrak{g},\mathfrak{g}]$
is a proper ideal of $\mathfrak{g}$. This is an open question and has been explicitly conjectured in [Gor1, Gor2]. The difficulty in addressing this problem reflects the more rigid nature of hypercomplex geometry compared to the complex case.
A central geometric object in this setting is the Obata connection: the unique torsion-free connection preserving all three complex structures. Its holonomy group $\operatorname{Hol}(\nabla)$ lies in $\operatorname{GL}(n,\mathbb{H})$ and plays a crucial role in understanding the geometry of hypercomplex manifolds [IP, SV]. In general, explicit descriptions of $\operatorname{Hol}(\nabla)$ are rare (see, for instance, [Sol, AGT, BFGV]). For nilmanifolds, however, it is known that the holonomy is contained in the commutator subgroup $\operatorname{SL}(n,\mathbb{H})$ [BDV, Corollary 3.3].
In this article, we investigate left-invariant hypercomplex structures on $2$-step nilpotent Lie groups. Surprisingly, all known examples in this class are Obata-flat.
This naturally leads to the question: Are all 2-step nilpotent hypercomplex Lie algebras Obata-flat? We answer this in the negative, and clarify the role of the nilpotency step of each complex structure (see Sections 2 and 3 for details). In particular, we show that if the Obata connection is not flat, then each $J_{\alpha}$ must be 3-step nilpotent. No such examples had been previously known.
In Section 4, we construct examples of 2-step nilpotent hypercomplex Lie algebras in which one, two, or all three complex structures are 3-step nilpotent. Among these, Example 4.3 provides the first known example of a 2-step nilpotent hypercomplex Lie algebra with non-flat Obata connection. We also provide new examples of $k$-step nilpotent hypercomplex nilmanifolds, with arbitrary $k$, which are not Obata flat.
In Section 6, we prove that the holonomy algebra of the Obata connection is always abelian for $2$-step nilpotent hypercomplex Lie groups, and we describe it explicitly (Theorem 6.1). This reveals a surprising rigidity: even when the Obata connection is non-flat, the holonomy remains severely restricted. Our approach relies on proving the Gorginyan conjecture (Conjecture 5.3) in the 2-step case (Theorem 5.4). We further show that the Obata holonomy group of a $2$-step hypercomplex nilmanifold is trivial if and only if the nilmanifold is a torus.
In the final part of the paper, we extend our analysis to the $3$-step nilpotent case, showing that, under suitable conditions, the holonomy of the Obata connection remains abelian.
Acknowledgments. A. Andrada and M. L. Barberis were partially supported by CONICET, SECyT-UNC and ANPCyT (Argentina). B. Brienza was partially supported by Project PRIN 2022 “Real and complex manifolds: Geometry and Holomorphic Dynamics”, by GNSAGA (Indam) and by the INdAM - GNSAGA Project CUP E53C24001950001. The third author is grateful to A. Fino, G. Gentili, G. Gorginyan, and M. Verbitsky for many helpful discussions, and to the FAMAF for the warm hospitality during the preparation of this work.
2 Preliminaries
A complex structure on a differentiable manifold $M$ is an automorphism $J$ of the tangent bundle $TM$ satisfying $J^{2}=-I$ and the integrability condition $N_{J}(X,Y)=0$ for all vector fields $X,Y$ on $M$, where $N_{J}$ denotes the Nijenhuis tensor defined by
$N_{J}(X,Y)=[X,Y]+J([JX,Y]+[X,JY])-[JX,JY].$
The integrability of $J$ is equivalent to the existence of an atlas on $M$ whose transition functions are holomorphic maps [NN].
A hypercomplex structure on $M$ is given by a triple of complex structures $\{J_{\alpha}\}$, $\alpha=1,2,3$, satisfying the quaternionic relations
$J_{1}J_{2}=-J_{2}J_{1}=J_{3}.$ (2.1)
As a consequence, $M$ admits a 2-sphere family of complex structures $J_{y}=y_{1}J_{1}+y_{2}J_{2}+y_{3}J_{3},$ parametrized by points $y=(y_{1},y_{2},y_{3})$ in the unit sphere $S^{2}\subset\mathbb{R}^{3}$. It follows from (2.1) that at each point $p\in M$, the tangent space $T_{p}M$ carries a natural structure of a module over the quaternions $\mathbb{H}$. In particular, $\dim M\equiv 0\;\pmod{4}$.
Given a hypercomplex structure $\{J_{\alpha}\}$ on a manifold $M$, there exists a unique torsion-free connection $\nabla$ on $M$ that preserves each complex structure, i.e., satisfies $\nabla J_{\alpha}=0$ for $\alpha=1,2,3$. This connection, known as the Obata connection (see [Ob, Sol]), was shown in [Sol] to admit the following explicit expression:
$\nabla_{X}Y=\frac{1}{2}\bigl{(}[X,Y]+J_{1}[J_{1}X,Y]-J_{2}[X,J_{2}Y]+J_{3}[J_{1}X,J_{2}Y]\bigr{)}.$ (2.2)
More generally, by applying cyclic permutations $(\alpha,\beta,\gamma)$ of $(1,2,3)$, the Obata connection can also be written as:
$\nabla_{X}Y=\frac{1}{2}\bigl{(}[X,Y]+J_{\alpha}[J_{\alpha}X,Y]-J_{\beta}[X,J_{\beta}Y]+J_{\gamma}[J_{\alpha}X,J_{\beta}Y]\bigr{)}.$ (2.3)
It is known that $(M,\{J_{\alpha}\})$ is quaternionic, (i.e., locally modelled over $\mathbb{H}^{n}$) if and only if the Obata connection $\nabla$ is flat [Ob, Theorem 11.2]. This highlights a fundamental difference with the complex case. For instance, the $K3$ surfaces and the homogeneous spaces $\operatorname{SU}(3)$ and $\operatorname{SO}(6)/\operatorname{SU}(2)$ provide examples of hypercomplex but not quaternionic manifolds [Boy, Joy].
Let $G$ be a Lie group with Lie algebra $\mathfrak{g}$. A hypercomplex structure $\{J_{\alpha}\}$ on $G$ is said to be left invariant if left translations by elements of $G$ are hyperholomorphic, i.e., holomorphic with respect to each $J_{\alpha}$. In this case, the structure is completely determined by its value at the identity, which corresponds to a hypercomplex structure on $\mathfrak{g}$.
If $\Gamma$ is a lattice in $G$, any left invariant hypercomplex structure on $G$ descends to a hypercomplex structure on the compact quotient $\G\Gamma\backslash$ G, which is then called invariant. In this setting, the natural projection $\GG\to\Gamma\backslash G$ is hyperholomorphic.
We recall the following definition.
Definition 2.1.
A Lie algebra $\mathfrak{g}$ is said to be nilpotent if its lower central series
$\mathfrak{g}=\mathfrak{g}^{0}\supset\mathfrak{g}^{1}\supset\cdots\supset\mathfrak{g}^{k}\supset\cdots,$
defined recursively by $\mathfrak{g}^{k}=[\mathfrak{g},\mathfrak{g}^{k-1}]$ for $k\geq 1$, eventually becomes zero; that is, there exists some $k\in\mathbb{N}$ such that
$\mathfrak{g}^{k}=0.$
The smallest such integer $k$ is called the nilpotency step of $\mathfrak{g}$.
Accordingly, a Lie algebra $\mathfrak{g}$ is said to be 2-step nilpotent if its commutator ideal $\mathfrak{g}^{1}=[\mathfrak{g},\mathfrak{g}]$ lies in the center $\mathfrak{z}$ of $\mathfrak{g}$. A Lie group $N$ is 2-step nilpotent if its Lie algebra $\mathfrak{g}$ is 2-step nilpotent.
Definition 2.2.
A 2-step nilmanifold is a compact quotient $\N\Gamma\backslash$ N, where $N$ is a connected, simply connected 2-step nilpotent Lie group and $\Gamma\subset N$ is a lattice.
It is well known that the existence of a lattice for a connected and simply connected nilpotent Lie group $N$ is equivalent to the existence of a rational basis of its Lie algebra $\mathfrak{g}$ [Mal].
Let $(\mathfrak{g},J)$ be a nilpotent Lie algebra equipped with a complex structure $J$. In this context, it is natural to consider an ascending series adapted to $J$.
Definition 2.3 ([CFGU]).
The ascending series $\{\mathfrak{a}_{k}\ ;\ k\geq 0\}$ (compatible with $J$) of $\mathfrak{g}$ is defined inductively by
$\mathfrak{a}_{0}=0,\ \mathfrak{a}_{k}=\{X\in\mathfrak{g}\ |\ [X,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\ \ \text{and}\ \ [JX,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\},\ k\geq 1.$ (2.4)
The complex structure $J$ is said to be nilpotent if there exists a $k\in\mathbb{N}$ such that $\mathfrak{a}_{k}=\mathfrak{g}$. The smallest such $k$ is called the nilpotency step of $J$.
If $\mathfrak{g}$ is $2$-step nilpotent, then any complex structure $J$ is nilpotent [Rol, Proposition 3.3], and the nilpotency step of $J$ is either $2$ or $3$ [GZZ, Theorem 1.3]. We recall from [B3] the following lemma.
Lemma 2.4 ([B3, Lemmas 3.3 and 4.1]).
Let $\mathfrak{g}$ be a $2$-step nilpotent Lie algebra endowed with a complex structure $J$. Then
  1. $(1)$
    $J$ is $2$-step nilpotent if and only if $J\mathfrak{g}^{1}\subset\mathfrak{z}$,
  2. $(2)$
    $J$ is $3$-step nilpotent if and only if there exists $X\in\mathfrak{g}^{1}$ such that $JX\notin\mathfrak{z}$.
  3. $(3)$
    $J\mathfrak{g}^{1}$ is an abelian ideal of $\mathfrak{g}$.
  • Proof.
    We include the proof of Statement (3) for further purposes. The proof is an application of the integrability of $J$. For any $X\in\mathfrak{g}$ and $Y\in\mathfrak{g}^{1}$, $N_{J}=0$ implies that
    [JY,X]=J[JY,JX].[JY,X]=-J[JY,JX]. (2.5)
    In particular, if $X\in J\mathfrak{g}^{1}$, then $[JY,X]=0$, as $\mathfrak{g}^{1}\subset\mathfrak{z}$. ∎
3 Computation of the Obata curvature
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. As already observed, the commutator ideal $\mathfrak{g}^{1}=[\mathfrak{g},\mathfrak{g}]$ lies in the center $\mathfrak{z}$ of $\mathfrak{g}$. This special feature allows for an explicit computation of the curvature tensor of the Obata connection.
Proposition 3.1.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. Then, for any $X,Y,Z\in\mathfrak{g}$ the curvature tensor of the Obata connection is given by
$4R(X,Y)Z=$ $-J_{\beta}[X,J_{\beta}[Y,Z]]+J_{\gamma}[J_{\alpha}X,J_{\beta}[Y,Z]]$
$-J_{\gamma}[X,J_{\beta}[J_{\alpha}Y,Z]]-J_{\beta}[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,Z]]$
$-[X,J_{\beta}[Y,J_{\beta}Z]]-J_{\alpha}[J_{\alpha}X,J_{\beta}[Y,J_{\beta}Z]]$
$+J_{\alpha}[X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]]-[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]]$
$+J_{\beta}[Y,J_{\beta}[X,Z]]-J_{\gamma}[J_{\alpha}Y,J_{\beta}[X,Z]]$
$+J_{\gamma}[Y,J_{\beta}[J_{\alpha}X,Z]]+J_{\beta}[J_{\alpha}Y,J_{\beta}[J_{\alpha}X,Z]]$ (3.1)
$+[Y,J_{\beta}[X,J_{\beta}Z]]+J_{\alpha}[J_{\alpha}Y,J_{\beta}[X,J_{\beta}Z]]$
$-J_{\alpha}[Y,J_{\beta}[J_{\alpha}X,J_{\beta}Z]]+[J_{\alpha}Y,J_{\beta}[J_{\alpha}X,J_{\beta}Z]]$
$-2[J_{\beta}[X,Y],J_{\beta}Z]+2J_{\gamma}[J_{\beta}[X,Y],J_{\alpha}Z],$
where $(\alpha,\beta,\gamma)$ is any cyclic permutation of $(1,2,3)$.
  • Proof.
    The proof is an application of the formula (2.3) and the integrability of $\{J_{\alpha}\}$. Using the formula (2.3) twice, the inclusion $\mathfrak{g}^{1}\subset\mathfrak{z}$ and that $\nabla_{X}J_{\alpha}Y=J_{\alpha}\nabla_{X}Y$, we compute:
    $4\nabla_{X}\nabla_{Y}Z$ $=-J_{\beta}[X,J_{\beta}[Y,Z]]+J_{\gamma}[J_{\alpha}X,J_{\beta}[Y,Z]]$
    $\quad-J_{\gamma}[X,J_{\beta}[J_{\alpha}Y,Z]]-J_{\beta}[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,Z]]$
    $\quad-[X,J_{\beta}[Y,J_{\beta}Z]]-J_{\alpha}[J_{\alpha}X,J_{\beta}[Y,J_{\beta}Z]]$
    $\quad+J_{\alpha}[X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]]-[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]].$
    Similarly, applying again (2.3) and using the inclusion $\mathfrak{g}^{1}\subset\mathfrak{z}$, we compute:
    $2\nabla_{[X,Y]}Z$ $=J_{\alpha}[J_{\alpha}[X,Y],Z]+J_{\gamma}[J_{\alpha}[X,Y],J_{\beta}Z]$
    $=[J_{\alpha}[X,Y],J_{\alpha}Z]+[J_{\beta}[X,Y],J_{\beta}Z]$
    $\quad-[J_{\alpha}[X,Y],J_{\alpha}Z]-J_{\gamma}[J_{\beta}[X,Y],J_{\alpha}Z]$
    $=[J_{\beta}[X,Y],J_{\beta}Z]-J_{\gamma}[J_{\beta}[X,Y],J_{\alpha}Z],$
    where the second equality follows from the integrability of $J_{\alpha}$ on the first term and the integrability of $J_{\gamma}$ on the second term. ∎
As a consequence of Proposition 3.1 we obtain the following result. This is a generalization of [B1, Proposition 5.1], where it was shown that if $\mathfrak{z}$ is $J_{\alpha}$ invariant for all $\alpha$ then $\{J_{\alpha}\}$ is Obata flat.
Corollary 3.2.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. If any of the three complex structures $J_{1},J_{2},J_{3}$ is $2$-step nilpotent, then $(\mathfrak{g},\{J_{\alpha}\})$ is Obata flat.
  • Proof.
    If $J_{\beta}$ is $2$-step nilpotent, $J_{\beta}\mathfrak{g}^{1}\subset\mathfrak{z}$ (Lemma 2.4). The result then follows from Proposition 3.1, since each term contains an element of the form $\cdot,J_{\beta}\mathfrak{g}^{1}]$. ∎
Remark 3.3.
As a consequence of Corollary 3.2, we have that if the Obata connection associated to $(\mathfrak{g},\{J_{\alpha}\})$ is not flat, then $J_{\alpha}$ is $3$-step for all $\alpha$. Moreover, any complex structure $J_{y}$ in the hypercomplex 2-sphere is $3$-step.
4 Construction of non-flat $2$-step hypercomplex nilmanifolds
To the best of our knowledge, in all known examples of $2$-step nilpotent hypercomplex Lie algebras $(\mathfrak{g},\{J_{\alpha}\})$, each complex structure $J_{\alpha}$ is $2$-step [B2, DF1, DF2, BD, LW, FG]. In particular, every known example of a $2$-step nilpotent Lie group equipped with a left-invariant hypercomplex structure is Obata flat (see Corollary 3.2).
According to [DF2, Section 3], in dimension $8$ any nilpotent Lie algebra admitting a hypercomplex structure is $2$-step nilpotent, and, moreover, each complex structure $J_{\alpha}$ is $2$-step. This naturally raises the question of whether this behavior is general among $2$-step nilpotent hypercomplex Lie algebras, or whether there exist examples in which one, two, or all of the complex structures are $3$-step.
Pushing this further: if such examples where each $J_{\alpha}$ is $3$-step exist, are the corresponding $2$-step nilpotent hypercomplex Lie groups still Obata flat? Looking at Formula (3.1), this is not clear.
In this section, we address these questions.
The construction
Let $\mathfrak{n}$ be a $2$-step nilpotent Lie algebra. We consider $\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{r}$.
Let $\mu\in\bigwedge^{2}\mathfrak{n}^{*}\otimes\mathbb{R}^{r}$ be a $2$-form on $\mathfrak{n}$ with values in $\mathbb{R}^{r}$ such that $\mathfrak{n}^{1}\subset\ker(\mu)$. We use $\mu$ to define a bracket $\cdot$$\cdot$$^{\prime}$ on $\mathfrak{g}$ as follows:
$[X,Y]^{\prime}=([X,Y],\mu(X,Y)),\ X,Y\in\mathfrak{n},$
$[X,Y]^{\prime}=0,\ X,Y\in\mathbb{R}^{r},$
$[X,Y]^{\prime}=0,\ X\in\mathfrak{n}\ \ \text{and}\ \ Y\in\mathbb{R}^{r}.$
Then, $(\mathfrak{g},[\cdot,\cdot]^{\prime})$ is a $2$-step nilpotent Lie algebra.
Let us now assume that $\mathfrak{n}$ is a $2$-step nilpotent Lie algebra endowed with a hypercomplex structure $\{J_{\alpha}\}$, and let $(\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k},[\cdot,\cdot]^{\prime})$ be the $2$-step nilpotent Lie algebra constructed as above. $\mathfrak{g}$ carries a natural almost complex structure $J_{\alpha}^{\prime}=J_{\alpha}\oplus I_{\alpha}$, where $\{I_{\alpha}\}$ is the standard hypercomplex structure on $\mathbb{R}^{4k}$. The following result is a characterization of $N_{J^{\prime}_{\alpha}}=0$.
Lemma 4.1.
The almost complex structure $J^{\prime}_{\alpha}$ is integrable if and only if
$I_{\alpha}\mu(X,Y)=\mu(J_{\alpha}X,Y)+\mu(X,J_{\alpha}Y)+I_{\alpha}\mu(J_{\alpha}X,J_{\alpha}Y),\ X,Y\in\mathfrak{n}.$ (4.1)
Let us now assume that the $2$-step nilpotent hypercomplex Lie algebra $(\mathfrak{n},\{J_{\alpha}\})$ satisfies that each $J_{\alpha}$ is $2$-step. If there exists $X\in\mathfrak{n}^{1}$ such that $\mu(J_{\alpha}X,\cdot)\neq 0$, then the corresponding complex structure $J^{\prime}_{\alpha}$ on $\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k}$ is $3$-step, according to Lemma 2.4. We will use this procedure to construct new examples of $2$-step nilpotent hypercomplex Lie algebras where one, two or all three complex structures are $3$-step.
Examples
Let $\mathfrak{n}_{8}$ be the $8$-dimensional 22-step nilpotent Lie algebra
$de^{i}=0,\ \ i=1,\dots,7,\ \ de^{8}=e^{12}-e^{34}.$ (4.2)
It has been shown in [DF1, DF2] that $\mathfrak{n}_{8}$ admits a hypercomplex structure $\{J_{\alpha}\}$ defined as
$J_{1}e_{1}=e_{2},\ J_{1}e_{3}=e_{4},\ J_{1}e_{5}=e_{6},\ J_{1}e_{7}=e_{8},$
$J_{2}e_{1}=e_{3},\ J_{2}e_{2}=-e_{4},\ J_{2}e_{5}=e_{7},\ J_{2}e_{6}=-e_{8},$ (4.3)
$J_{3}e_{1}=e_{4},\ J_{3}e_{2}=e_{3},\ J_{3}e_{5}=e_{8},\ J_{3}e_{6}=e_{7}.$
In particular, since $\mathfrak{n}_{8}^{1}=\text{span}\{e_{8}\}$ and the center $\mathfrak{z}(\mathfrak{n}_{8})=\text{span}\{e_{5},e_{6},e_{7},e_{8}\}$, $J_{\alpha}$ is $2$-step for any $\alpha=1,2,3$.
Consider $\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4}$ and fix the standard basis $\{e_{9},e_{10},e_{11},e_{12}\}$ of $\mathbb{R}^{4}$. We point out that $\{I_{\alpha}\}$ are defined as
$I_{1}e_{9}=e_{10},\ I_{1}e_{11}=e_{12},$
$I_{2}e_{9}=e_{11},\ I_{2}e_{10}=-e_{12},$
$I_{3}e_{9}=e_{12},\ I_{3}e_{10}=e_{11}.$
Example 4.2.
We consider the following Lie bracket $\cdot$$\cdot$$^{\prime}$ on $\mathfrak{g}$ given by $\cdot,\cdot]^{\prime}=[\cdot,\cdot]+\mu$, where $\mu\in\bigwedge^{2}\mathfrak{n}_{8}^{*}\otimes\mathbb{R}^{4}$ is defined as
$\mu=e^{15}\otimes e_{9}+e^{25}\otimes e_{10}+e^{35}\otimes e_{11}+e^{45}\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. In particular, since $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{3}$ is $3$-step nilpotent, while $J^{\prime}_{1},J^{\prime}_{2}$ are $2$-step nilpotent.
To sum up, the $2$-step nilpotent Lie algebra $\mathfrak{g}$ defined by the structure equations
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{12}-e^{34},\\ &de^{9}=-e^{15},\ \ de^{10}=-e^{25},\\ &de^{11}=-e^{35},\ \ de^{12}=-e^{45},\end{split}$
can be endowed with a hypercomplex structure $\{J^{\prime}_{\alpha}\}$, where $J^{\prime}_{1},J^{\prime}_{2}$ are $2$-step nilpotent and $J^{\prime}_{3}$ is $3$-step nilpotent.
In order to construct a $2$-step nilpotent hypercomplex Lie algebra in which two of the complex structures are $3$-step nilpotent and one is $2$-step nilpotent, we consider the same underlying vector space $\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4}$ as above, but we modify the Lie bracket by changing the $2$-form $\mu$. Let us define on $\mathfrak{g}$ the Lie bracket $\cdot,\cdot]^{\prime}=[\cdot,\cdot]+\mu$, where $\mu\in\bigwedge^{2}\mathfrak{n}_{8}^{*}\otimes\mathbb{R}^{4}$ is now defined as
$\mu=\big{(}e^{15}+e^{16}\big{)}\otimes e_{9}+\big{(}e^{25}+e^{26}\big{)}\otimes e_{10}+\big{(}e^{35}+e^{36}\big{)}\otimes e_{11}+\big{(}e^{45}+e^{46}\big{)}\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. In particular, since $\mu(J_{2}e_{8},\cdot)=\mu(e_{6},\cdot)\neq 0$ and $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent, while $J^{\prime}_{1}$ is $2$-step nilpotent.
In summary, the $2$-step nilpotent Lie algebra $\mathfrak{g}$ defined by the structure equations
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ \ de^{8}=e^{12}-e^{34},\\ &de^{9}=-e^{15}-e^{16},\ \ de^{10}=-e^{25}-e^{26},\\ &de^{11}=-e^{35}-e^{36},\ \ de^{12}=-e^{45}-e^{46},\end{split}$
can be endowed with a hypercomplex structure $\{J^{\prime}_{\alpha}\}$, where $J^{\prime}_{1}$ is $2$-step nilpotent and $J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent.
We point out that both of examples constructed above are still Obata flat by Corollary 3.2.
Example 4.3.
In this last example we construct a $2$-step nilpotent hypercomplex Lie algebra where each complex structure is $3$-step. Let us consider again $\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4}$ with bracket $\cdot$$\cdot$$^{\prime}$ defined as $\cdot,\cdot]^{\prime}=[\cdot,\cdot]+\mu$ where $\mu\in\bigwedge^{2}\mathfrak{n}_{8}^{*}\otimes\mathbb{R}^{4}$ has the following definition:
$\mu=e^{56}\otimes e_{9}-e^{67}\otimes e_{10}+e^{57}\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. Since $\mu(J_{1}e_{8},\cdot)=\mu(-e_{7},\cdot)\neq 0$$\mu(J_{2}e_{8},\cdot)=\mu(e_{6},\cdot)\neq 0$ and $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{1},J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent complex structures.
To sum up, the $2$-step nilpotent Lie algebra $\mathfrak{g}$ defined by the structure equations
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{12}-e^{34},\\ &de^{9}=-e^{56},\ \ de^{10}=0,\ \ \ de^{11}=e^{67},\ \ de^{12}=-e^{57},\end{split}$
admits a hypercomplex structure $\{J^{\prime}_{\alpha}\}$, where $J^{\prime}_{1},J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent. We claim that this hypercomplex Lie algebra is not Obata flat. Using Formula (2.2):
$\begin{split}R^{\prime}(e_{8},e_{1})e_{1}&=\nabla^{\prime}_{e_{8}}\nabla^{\prime}_{e_{1}}e_{1}=-\frac{1}{2}\nabla^{\prime}_{e_{8}}e_{7}=-\frac{1}{4}e_{9},\end{split}$
where we used that $\nabla^{\prime}_{e_{8}}e_{1}=0$ and $e_{1},e_{8}]=0$.
In Section 6, we will prove that if we consider the corresponding $2$-step nilpotent hypercomplex Lie group, then the holonomy algebra $\mathfrak{hol}(\nabla^{\prime})$ is abelian.
We point out that the $2$-step nilpotent Lie groups corresponding to the three Lie algebras constructed in this section admit lattices [Mal].
Semidirect product construction
We present next a construction which is a slight generalization of [BF, Lemma 3.1].
Let $(\mathfrak{g},\{J_{\alpha}\})$ be an arbitrary hypercomplex Lie algebra and let $\{I_{\alpha}\}$ be the standard hypercomplex structure on $\mathbb{R}^{4k}$. Let $\rho:\mathfrak{g}\to\mathfrak{g}\mathfrak{l}(k,\mathbb{H})$ be a Lie algebra homomorphism, that is,
$\rho([X,Y])=[\rho(X),\rho(Y)],\quad\rho(X)I_{\alpha}={I_{\alpha}}_{\,}\rho(X),\;\;\text{ for all }X,\,Y\in\mathfrak{g},\;\alpha=1,2,3.$
Consider the Lie algebra $\mathfrak{h}:=\mathfrak{g}\ltimes_{\rho}\mathbb{R}^{4k}$, that is,
${[(X,V),(Y,W)]}=([X,Y],\rho(X)W-\rho(Y)V),\quad X,Y\in\mathfrak{g},\;V,W\in\mathbb{R}^{4k}.$
The Jacobi identity on $\mathfrak{h}$ follows from the fact that $\rho$ is a Lie algebra homomorphism. Note that $\{(0,V):V\in\mathbb{R}^{4k}\}$ is an abelian ideal of $\mathfrak{h}$. Let $\widetilde{J}_{\alpha}$, $\alpha=1,2,3$, be the endomorphisms of $\mathfrak{h}$ defined by $\widetilde{J}_{\alpha}=J_{\alpha}\oplus I_{\alpha}$.
Proposition 4.4.
Let $(\mathfrak{g},\{J_{\alpha}\})$ and $(\mathfrak{h},\{\widetilde{J}_{\alpha}\})$ be as above, then $\{\widetilde{J}_{\alpha}\}$ is a hypercomplex structure on $\mathfrak{h}$. Moreover, the Obata connection $\widetilde{\nabla}$ of $\{\widetilde{J}_{\alpha}\}$ and the curvature tensor $\widetilde{R}$ of $\widetilde{\nabla}$ are given by:
$\widetilde{\nabla}_{(X,V)}(Y,W)$ $=$ $(\nabla_{X}Y,\rho(X)W),$
$\widetilde{R}\left((X,V),(Y,W)\right)(Z,U)$ $=$ $\left(R(X,Y)Z,0\right),\quad X,Y,Z\in\mathfrak{g},\;V,W,U\in\mathbb{R}^{4k},$
where $\nabla$ is the Obata connection of $\{J_{\alpha}\}$ and $R$ is the curvature tensor of $\nabla$. In particular, $\mathfrak{hol}(\widetilde{\nabla})\cong\mathfrak{hol}(\nabla)$.
  • Proof.
    Since $\{J_{\alpha}\}$ is integrable on $\mathfrak{g}$ and $\rho(X)$ commutes with $I_{\alpha}$ for all $X\in\mathfrak{g}$, $\alpha=1,2,3$, it turns out that $\{\widetilde{J}_{\alpha}\}$ is integrable on $\mathfrak{h}$.
    The rest of the proof follows the lines of [BF, Lemma 3.1]. ∎
Lemma 4.5.
  1. $(1)$
    The center and commutator ideal of the Lie algebra $\mathfrak{h}$ are given by:
    $\mathfrak{z}(\mathfrak{h})$ $=$ $\left\{(X,V):X\in\mathfrak{z}(\mathfrak{g})\cap\ker\rho,\;V\in\bigcap_{\begin{subarray}{c}Y\in\mathfrak{g}\end{subarray}}\ker\rho(Y)\right\},$
    $\mathfrak{h}^{1}$ $=$ $\left\{(X,V):X\in\mathfrak{g}^{1},\;V\in\rho(\mathfrak{g})\mathbb{R}^{4k}\right\}.$
  2. $(2)$
    If $\mathfrak{g}$ is $l$-step nilpotent and
    $m_{\rho}=\min\{j:\rho(X_{1})\cdots\rho(X_{j})=0,\ \text{for all}\ \ X_{1},\dots,X_{j}\in\mathfrak{g}\},$
    then $\mathfrak{h}$ is $k$-step nilpotent, where $k=\max\{l,m_{\rho}\}$. In particular, if $\mathfrak{g}$ is $2$-step nilpotent and $\rho(X_{1})\cdot\rho(X_{2})=0$ for any $X_{1},X_{2}\in\mathfrak{g}$, then $\mathfrak{h}$ is $2$-step nilpotent.
  3. $(3)$
    If $\mathfrak{h}$ is 2-step nilpotent, then $\widetilde{J}_{\alpha}$ is 2-step nilpotent if and only if $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g})\cap\ker\rho$. In particular, if $J_{\alpha}$ is 3-step nilpotent, then $\widetilde{J}_{\alpha}$ is 3-step nilpotent.
  • Proof.
    The only non-trivial statement is the second, which follows from the identity
    $\operatorname{ad}_{(X,U)}^{k}=\left(\begin{array}[]{c|c}\operatorname{ad}_{X}^{k}&0\\ \hline\cr B_{U}^{k}&\rho(X)^{k}\end{array}\right),$
    where the operator $B_{U}^{k}:\mathfrak{g}\to\mathbb{R}^{4k}$ is given by
    $B_{U}^{k}Y=\left(\sum_{l=1}^{k}(-1)^{l}\binom{k}{l}\rho(X)^{k-l}\rho(Y)\rho(X)^{l-1}\right)U.$
New examples of Obata non flat 2-step nilpotent hypercomplex Lie algebras can be obtained by applying the semidirect product construction, as we show next.
Example 4.6.
Let $\mathfrak{g}$ be the Lie algebra from Example 4.3 with the hypercomplex structure considered there. We define $\rho:\mathfrak{g}\to\mathfrak{g}\mathfrak{l}(2,\mathbb{H})\subset\mathfrak{gl}(8,\mathbb{R})$ by:
$\rho(e_{1})=\begin{bmatrix}0_{4}&0_{4}\\ I_{4}&0_{4}\end{bmatrix},\qquad\quad\rho(e_{i})=0,\quad i\geq 2,$
[…]
Example 4.7.
Let $\mathfrak{g}$ be as in Example 4.3 and set $\rho:\mathfrak{g}\to\mathfrak{g}\mathfrak{l}(k,\mathbb{H})\subset\mathfrak{gl}(4k,\mathbb{R})$, $k\geq 3$, defined by:
$\rho(e_{1})=\left[\begin{array}[]{ccccc}0_{4}&0_{4}&\cdots&\cdots&0_{4}\\ I_{4}&0_{4}&0_{4}&\cdots&0_{4}\\ 0_{4}&I_{4}&0_{4}&\cdots&0_{4}\\ \vdots&\vdots&\ddots&\ddots&0_{4}\\ 0_{4}&0_{4}&0_{4}&I_{4}&0_{4}\end{array}\right]\in M_{4k}(\mathbb{R}),\qquad\rho(e_{i})=0,\quad i\geq 2.$
In this case, $\mathfrak{h}$ is $k$-step nilpotent and $\{\widetilde{J}_{\alpha}\}$ is not Obata flat. By Proposition 4.4, the holonomy algebra coincides with the holonomy algebra of the Example 4.3. In particular, it is abelian (see Section 6).
We note that the simply connected Lie groups corresponding to the Lie algebras from examples 4.6 and 4.7 admit lattices, according to [Mal].
5 $\mathbb{H}$-solvability of $2$-step nilpotent Lie algebras
We recall the following definition.
Definition 5.1 ([Gor1]).
Let $\mathfrak{g}$ be a nilpotent Lie algebra endowed with a hypercomplex structure $(J_{1},J_{2},J_{3})$. Define inductively $\mathbb{H}$-invariant Lie subalgebras:
$\mathfrak{g}_{k}^{\mathbb{H}}:=\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]+J_{1}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]+J_{2}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]+J_{3}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}],$
where $\mathfrak{g}_{0}^{\mathbb{H}}=\mathfrak{g}$. Note that $\mathfrak{g}_{1}^{\mathbb{H}}:=\mathbb{H}\mathfrak{g}^{1}=\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1}$.
A hypercomplex nilpotent Lie algebra $\mathfrak{g}$ is called $\mathbb{H}$-solvable if the following sequence eventually becomes zero for some $k\in\mathbb{N}$:
$\mathfrak{g}_{1}^{\mathbb{H}}\supset\mathfrak{g}_{2}^{\mathbb{H}}\supset\cdots\supset\mathfrak{g}_{k-1}^{\mathbb{H}}\supset\mathfrak{g}_{k}^{\mathbb{H}}=0.$ (5.1)
The smallest such $k$ is called the $\mathbb{H}$-solvability step of $(\mathfrak{g},\{J_{\alpha}\})$.
Lemma 5.2.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a nilpotent hypercomplex Lie algebra. Then for any $k\in\mathbb{N}$, $\mathfrak{g}_{k}^{\mathbb{H}}$ is a subalgebra of $\mathfrak{g}$ and it is an ideal in $\mathfrak{g}_{k-1}^{\mathbb{H}}$.
  • Proof.
    Since $\mathfrak{g}^{1}\subset\mathfrak{g}^{1}_{\mathbb{H}}$, $\mathfrak{g}^{\mathbb{H}}_{1}$ is clearly a subalgebra. By induction, assume that $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is a subalgebra. Then
    $\mathfrak{g}_{k}^{\mathbb{H}},\mathfrak{g}_{k}^{\mathbb{H}}]=[\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}],\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]].$
    Since $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is a subalgebra $\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset\mathfrak{g}_{k-1}^{\mathbb{H}}$. In particular
    $\mathfrak{g}_{k}^{\mathbb{H}},\mathfrak{g}_{k}^{\mathbb{H}}]\subset[\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}},\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset\mathfrak{g}^{\mathbb{H}}_{k}.$
    Since $\mathfrak{g}_{k}^{\mathbb{H}}$ is a subalgebra for each $k$, we have:
    $[\mathfrak{g}_{k}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}],\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset[\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset\mathfrak{g}_{k}^{\mathbb{H}}.$
Let $(\mathfrak{g},J)$ be a nilpotent Lie algebra endowed with a complex structure $J$. It was shown by Salamon [Sal, Corollary 1.4] that the ideal $\mathfrak{g}^{1}+J\mathfrak{g}^{1}$ is a proper ideal of $\mathfrak{g}$. This naturally raises the question of whether an analogous property holds for nilpotent Lie algebras endowed with a hypercomplex structure.
The condition that $\mathfrak{g}_{1}^{\mathbb{H}}=\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1}$ is a proper ideal of $\mathfrak{g}$ is necessary to the $\mathbb{H}$-solvability of $(\mathfrak{g},\{J_{\alpha}\})$. This observation motivates the following conjecture:
Conjecture 5.3 ([Gor1]).
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a nilpotent hypercomplex Lie algebra. Then it is $\mathbb{H}$-solvable.
When the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ is $\mathbb{H}$-solvable, it turns out that, for a generic complex structure $J_{y}$ in the hypercomplex $2$-sphere, the complex nilmanifold $(\Gamma\backslash G,J_{y})$ does not admit any complex curves [Gor1].
The conjecture holds true for Lie algebras of hypercomplex nilmanifolds with flat Obata connection [Gor2] and for nilpotent Lie algebras with abelian hypercomplex structure [AV, Proposition 4.5]. Here we prove the conjecture for $2$-step nilpotent Lie algebras.
Theorem 5.4.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. Then $(\mathfrak{g},\{J_{\alpha}\})$ is $\mathbb{H}$-solvable. In particular, $\mathfrak{g}_{1}^{\mathbb{H}}$ is a proper ideal of $\mathfrak{g}$ and the $\mathbb{H}$-solvability step is at most $3$.
  • Proof.
    To prove Theorem 5.4, we show that the $\mathbb{H}$-solvable series (5.1) terminates at 0. In particular, this implies that for any $k\in\mathbb{N}_{\geq 1}$, $\mathfrak{g}_{k}^{\mathbb{H}}\subsetneq\mathfrak{g}$, proving that $\mathfrak{g}_{1}^{\mathbb{H}}$ is a proper ideal of the Lie algebra $\mathfrak{g}$.
    Let us consider
    $\mathfrak{g}_{2}^{\mathbb{H}}=\mathbb{H}[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}].$
    By Lemma 2.4,
    $[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}]=$ $[\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1},\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1}]$
    $=$ $[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}].$
    Therefore,
    $\mathbb{H}[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}]=$ $\ [J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]$
    $+J_{1}([J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}])$
    $+J_{2}([J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}])$
    $+J_{3}([J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]).$
    Since by Equation 2.5 $J_{\alpha}[J_{\alpha}\mathfrak{g}^{1},J_{\beta}\mathfrak{g}^{1}]\subset\mathfrak{g}^{1}$, we may write:
    $\mathfrak{g}^{\mathbb{H}}_{2}=\mathbb{H}[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}]$ $\subset\mathfrak{g}^{1}+J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}].$
    We claim that $J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]$. Using the quaternionic identities and Equation 2.5 with $J_{1}$, we have:
    $J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}J_{1}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}],$
    and, similarly, using the integrability of $J_{2}$,
    $J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]=J_{1}J_{2}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]=J_{1}[J_{3}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}].$
    Thus, all three summands coincide and we obtain that $\mathfrak{g}^{\mathbb{H}}_{2}\subset\mathfrak{g}^{1}+J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]\subset\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}$. Therefore,
    $\mathfrak{g}_{3}^{\mathbb{H}}=\mathbb{H}[\mathfrak{g}^{\mathbb{H}}_{2},\mathfrak{g}^{\mathbb{H}}_{2}]\subset\mathbb{H}[J_{1}\mathfrak{g}^{1},J_{1}\mathfrak{g}^{1}]=0,$
    since $J_{1}\mathfrak{g}^{1}$ is abelian by Lemma 2.4. ∎
6 Holonomy of $2$-step hypercomplex nilmanifolds
We use Theorem 5.4 to prove a structural result about the holonomy algebra of the Obata connection on $2$-step nilpotent Lie groups endowed with invariant hypercomplex structures.
Theorem 6.1.
Let $G^{4n}$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. Then the holonomy Lie algebra $\mathfrak{hol}(\nabla)$ is abelian, where $\nabla$ is the Obata connection of $(G,\{J_{\alpha}\})$.
In particular, it is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$ with trivial product, that is, for any $A,B\in\mathfrak{hol}(\nabla),\ A\cdot B=0$.
  • Proof.
    The proof is divided into $4$ steps. We start by proving
    • Step 1: for any $X,Y,Z\in\mathfrak{g}$, $R(X,Y)Z\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$.
      For $(\alpha,\beta,\gamma)=(1,2,3)$, each term appearing in Formula (3.1) falls into one of four distinct types:
      1. 1.
        $J_{2}[\cdot,J_{2}[\cdot,\cdot]]=[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, where the equality follows by Equation (2.5),
      2. 2.
        $\cdot,J_{2}[\cdot,\cdot]]\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, as $J_{2}\mathfrak{g}^{1}$ is an ideal by Lemma 2.4,
      3. 3.
        $J_{3}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows by Equation (2.5),
      4. 4.
        $J_{1}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows again by Equation (2.5).
      We set $\mathfrak{h}:=\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. It is not difficult to note that $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant and $\mathfrak{h}\subset\mathfrak{g}_{1}^{\mathbb{H}}$.
    • Step 2: for any $X\in\mathfrak{h}$, $\nabla X=0$.
      We use the following:
      Lemma 6.2.
      For any $X\in\mathfrak{h}$ and $Y\in\mathfrak{g}$,
      $Y,X]+J_{1}[J_{1}Y,X]=0$ (6.1)
      • Proof.
        Let $X\in\mathfrak{h}$. Then $X=X_{1}+X_{2}$, where $X_{1}\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$ and $X_{2}\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. Since $X_{1}\in\mathfrak{g}^{1}$,
        $Y,X]+J_{1}[J_{1}Y,X]=[Y,X_{2}]+J_{1}[J_{1}Y,X_{2}].$
        Using that $X_{2}\in J_{1}\mathfrak{g}^{1}$, $X_{2},Y]=-J_{1}[X_{2},J_{1}Y]$ by Equation (2.5). Therefore,
        $Y,X]+J_{1}[J_{1}Y,X]=-J_{1}[J_{1}Y,X_{2}]+J_{1}[J_{1}Y,X_{2}]=0.$
      Let $Y\in\mathfrak{g}$ and $X\in\mathfrak{h}$. Applying Formula (2.2)
      $2\nabla_{Y}X$ $=[Y,X]+J_{1}[J_{1}Y,X]-J_{2}[Y,J_{2}X]+J_{3}[J_{1}Y,J_{2}X]$
      $=[Y,X]+J_{1}[J_{1}Y,X]-J_{2}([Y,J_{2}X]+J_{1}[J_{1}Y,J_{2}X])$
      $=0,$
      where we used twice Lemma 6.2 and the $\{J_{\alpha}\}$-invariance of $\mathfrak{h}$.
    • Step 3: Expression for the curvature endomorphisms and $\nabla$.
      By Theorem 5.4, the ideal $\mathfrak{g}_{1}^{\mathbb{H}}$ is a proper ideal of $\mathfrak{g}$ containing $\mathfrak{h}$. Therefore, we may write $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{g}_{1}^{\mathbb{H}}$, where $\mathfrak{k}_{1}\neq 0$ is any $\{J_{\alpha}\}$-invariant complement of $\mathfrak{g}_{1}^{\mathbb{H}}$. Similarly, we decompose $\mathfrak{g}_{1}^{\mathbb{H}}=\mathfrak{h}_{1}\oplus\mathfrak{h}$, where $\mathfrak{h}_{1}$ is any $\{J_{\alpha}\}$-invariant complement of $\mathfrak{h}$ inside $\mathfrak{g}_{1}^{\mathbb{H}}$ (such complements always exist, as we can consider the orthogonal complements with respect to a fixed hyperHermitian metric). We note that $\mathfrak{h}_{1}$ could be trivial; nonetheless, the proof proceeds in exactly the same manner without any differences.
      With respect to the decomposition $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{h}_{1}\oplus\mathfrak{h}$, using Step $1$ and Step $2$, we get that for any $X,Y\in\mathfrak{g}$,
      $R(X,Y)=\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}$.
      We have the following two Lemmas:
      Lemma 6.3.
      Let $X,Y\in\mathfrak{g}$. Then $\nabla_{X}Y\in\mathfrak{g}_{1}^{\mathbb{H}}$.
      • Proof.
        Straightforward by Formula (2.2). ∎
      Lemma 6.4.
      Let $X\in\mathfrak{g}$. Then $\nabla_{X}\mathfrak{g}_{1}^{\mathbb{H}}\subset\mathfrak{h}$.
      • Proof.
        Since $\nabla_{X}J_{\alpha}=J_{\alpha}\nabla_{X}$ and $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant, it suffices to prove that $\nabla_{X}\mathfrak{g}^{1}\subset\mathfrak{h}$. Applying Formula (2.2), we see that for any $Y\in\mathfrak{g}^{1}$:
        $\nabla_{X}Y=\frac{1}{2}\big{(}-J_{2}[X,J_{2}Y]+J_{3}[J_{1}X,J_{2}Y]\big{)}.$
        By Step 1, the first term lies in $\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, while the second term lies in $J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. The proof follows by the definition of $\mathfrak{h}$. ∎
      By the previous Lemmas we get that with respect to the decomposition $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{h}_{1}\oplus\mathfrak{h}$:
      $\nabla_{X}=\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr\ast&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}$. (6.2)
    • Step 4: Application of the Ambrose-Singer Theorem.
      We recall the following version of the Ambrose-Singer Theorem for left invariant connections.
      Theorem 6.5.
      [Al] Let $\nabla$ be a left invariant linear connection on the Lie group $G$, and let $\mathfrak{g}$ denote the Lie algebra of $G$. Then the holonomy algebra $\mathfrak{hol}(\nabla)$, based at the identity element $e\in G$, is the smallest subalgebra of $\mathfrak{gl}(\mathfrak{g})$ containing the curvature endomorphisms $R(X,Y)$ for any $X,Y\in\mathfrak{g}$, and closed under commutators with the left multiplication operators $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$.
      Let us consider $\text{span}\{R(X,Y)\ |\ X,Y\in\mathfrak{g}\}$. As already observed in Step $3$, with respect to the decomposition $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{h}_{1}\oplus\mathfrak{h}$,
      $\text{span}\{R(X,Y)\ |\ X,Y\in\mathfrak{g}\}\subset\left\{\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}\in\mathfrak{sl}(n,\mathbb{H})\right\}.$
      We set $\mathfrak{a}:=\left\{\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}\in\mathfrak{sl}(n,\mathbb{H})\right\}$. We point out that $\mathfrak{a}$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$.
      Let $A\in\mathfrak{a}$ and let $X\in\mathfrak{g}$. Using the expression of $\nabla_{X}$ given in Step 3, the commutator
      $A,\nabla_{X}]\in\mathfrak{a}.$
      Since $\mathfrak{a}$ is closed under commutators with $\nabla_{\mathfrak{g}}$, by Theorem 6.5 we get that $\mathfrak{hol}(\nabla)\subset\mathfrak{a}$. In particular, since $\mathfrak{a}$ is abelian with trivial product, Theorem 6.1 is proved.
Example 6.6.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be the $2$-step nilmanifold hypercomplex Lie algebra constructed in Example 4.3 and let us consider the corresponding $2$-step nilpotent hypercomplex Lie group $G$. As already observed, the Obata connection $\nabla$ on $G$ is not flat. A direct computation shows that the holonomy algebra $\mathfrak{hol}(\nabla)$ is given by
$\mathfrak{hol}(\nabla)=\left\{\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr q&p\cdot j&0\end{array}\end{pmatrix}\in\mathfrak{sl}(3,\mathbb{H})\ |\ q\in\mathbb{H},\ p\in\mathbb{R}\right\},$
where we are considering $\mathfrak{g}=\text{span}_{\mathbb{H}}\{e_{1},e_{5},e_{9}\}$.
This provides an explicit example of a non-trivial, abelian holonomy algebra arising from a non-flat Obata connection on a $2$-step nilmanifold.
Let $G$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. If the Obata connection is flat, then the full holonomy group is discrete. A natural question is whether this could be trivial. To address this question, we prove the following theorem.
Theorem 6.7.
Let $(\Gamma\backslash G,\{J_{\alpha}\})$ be a $2$-step nilmanifold endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$ such that the Obata connection is flat. Then the holonomy group $\operatorname{Hol}(\nabla)$ is trivial if and only if $\G\Gamma\backslash$ G is a torus.
  • Proof.
    The implication from right to left is clear. Let us now prove the converse.
    The Obata connection is complex with respect to each $J_{\alpha}$ and has vanishing (and hence parallel) torsion. Moreover, as shown in the proof of Theorem 6.1, for any $X\in\mathfrak{g}$, the endomorphism $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$ is nilpotent (see Equation (6.2)). Since the Obata connection $\nabla$ is flat and torsion-free, it follows from [Kim, Theorem 2.2 and Proposition 1.2] that $\nabla$ is complete.
    We apply [ABD, Theorem 4.1] to conclude that $(\Gamma\backslash G,J_{\alpha},\nabla)$ is affinely biholomorphic to $(\Gamma_{0}\backslash G_{0},J_{\alpha}^{0},\nabla^{0})$ for each $\alpha$, where $G_{0}$ is the connected and simply connected Lie group whose Lie algebra is given by
[…]
    $\Gamma_{0}$ is a lattice in $G_{0}$, $J_{\alpha}^{0}$ is the complex structure induced on $P^{\nabla}$ by $J_{\alpha}$ and $\nabla^{0}$ is the ()(-)-connection.
[…]
Remark 6.8.
Theorem 6.7 still holds true if instead of considering a 2-step nilmanifold we assume that $\G\Gamma\backslash$ G is a solvmanifold with $G$ a completely solvable almost abelian Lie group. Indeed, in this case the Obata connection is automatically flat, according to [AB1, Proposition 3.7], and is complete, due to [AB1, Corollary 3.8]. For the last paragraph, we use the Saito rigidity Theorem [Sai], which extends the result of Mal’cev from nilpotent to completely solvable Lie groups.
7 Finale
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a Lie algebra endowed with a hypercomplex structure such that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for all $\alpha=1,2,3$. We claim that $\mathfrak{g}$ is at most $3$-step nilpotent. Indeed, using the integrability of $J_{\alpha}$ and the assumption $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, one obtains
$X,Y],Z]=-J_{\alpha}[[X,Y],J_{\alpha}Z],$ (7.1)
which shows that $\mathfrak{g}^{2}\subset\mathfrak{z}$, and hence $\mathfrak{g}^{3}=[\mathfrak{g}^{2},\mathfrak{g}]=0$. Note that by Equation (7.1), $\mathfrak{g}^{1},\mathfrak{g}^{1}]=0$.
Since in the $2$-step case the corresponding Obata connection is flat (see Corollary 3.2), we will focus our attention on the $3$-step case. We observe that the hypercomplex structure must be non-abelian. Indeed, if $\{J_{\alpha}\}$ is abelian and $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, then
$\mathfrak{g}^{1}\subset J_{\alpha}\mathfrak{z}=\mathfrak{z},$
which forces $\mathfrak{g}$ to be $2$-step nilpotent.
Examples of 3-step nilpotent Lie algebras with hypercomplex structures have been given in [DF3, BD, AB2]. Among these, the only examples with non-abelian hypercomplex structures are those in [AB2]; however, those are Obata flat and none of them satisfies the condition $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for all $\alpha=1,2,3$.
Due to the lack of examples of this kind in the literature, we include here a 3-step nilpotent example satisfying $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for any $\alpha=1,2,3$.
Let $\mathfrak{g}$ be the $3$-step nilpotent Lie algebra defined by the structure equations
$\begin{split}&de^{i}=0,\ i=1,\dots,4,\ \ de^{5}=e^{12}-e^{34},\\ &de^{i}=0,\ i=6,7,8,\ \ de^{9}=e^{13}-e^{42},\\ &de^{i}=0,\ i=10,11,12,\ \ de^{13}=-e^{25}+e^{39},\\ &de^{14}=e^{15}+e^{49},\ \ de^{15}=e^{45}-e^{19},\ \ de^{16}=-e^{35}-e^{29}.\end{split}$
Then $\mathfrak{g}$ can be endowed with the following hypercomplex structure
$J_{1}e_{1}=e_{2},\ J_{1}e_{3}=e_{4},\ J_{1}e_{5}=e_{6},\ J_{1}e_{7}=e_{8},$
$J_{1}e_{9}=e_{10},\ J_{1}e_{11}=e_{12},\,J_{1}e_{13}=e_{14},\ J_{1}e_{15}=e_{16},$
$J_{2}e_{1}=e_{3},\ J_{2}e_{2}=-e_{4},\ J_{2}e_{5}=e_{7},\ J_{2}e_{6}=-e_{8},$
$J_{2}e_{9}=e_{11},\ J_{2}e_{10}=-e_{12},J_{2}e_{13}=e_{15},\ J_{2}e_{14}=-e_{16},$
$J_{3}e_{1}=e_{4},\ J_{3}e_{2}=e_{3},\ J_{3}e_{5}=e_{8},\ J_{3}e_{6}=e_{7},$
$J_{3}e_{9}=e_{12},\ J_{3}e_{10}=e_{11},\ J_{3}e_{13}=e_{16},\ J_{3}e_{14}=e_{15},$
satisfying $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, for each $\alpha=1,2,3$.
Proposition 7.1.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a hypercomplex Lie algebra such that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for each $\alpha=1,2,3$. Then
$R(X,Y)=-\operatorname{ad}_{[X,Y]},$
and $(\mathfrak{g},\{J_{\alpha}\})$ is non-flat if and only if $\mathfrak{g}$ is $3$-step nilpotent.
Let $G$ be the corresponding connected and simply connected $3$-step nilpotent Lie group. Then the holonomy algebra of the Obata connection on $G$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$.
  • Proof.
    As we pointed out at the beginning of this section, $\mathfrak{g}$ is at most $3$-step nilpotent. Applying Formula (2.2) and the hypothesis that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for each $\alpha$, we get that
    $\nabla_{X}\nabla_{Y}Z$ $=\frac{1}{4}\left([X,[Y,Z]+J_{1}[J_{1}X,[Y,Z]]-J_{2}[X,[Y,J_{2}Z]]+J_{3}[J_{1}X,[Y,J_{2}Z]]\right)$
    $=0,$
    where we used Equation (7.1). Therefore,
    $R(X,Y)Z=-\nabla_{[X,Y]}Z=-\frac{1}{2}([[X,Y],Z]-J_{2}[[X,Y],J_{2}Z])=-[[X,Y],Z],$
    where the last equality again follows from Equation (7.1).
    Set $\mathfrak{a}:=\text{span}\{\operatorname{ad}_{[X,Y]}\ |\ X,Y\in\mathfrak{g}\}$. Then, $\mathfrak{a}$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$. In fact,
    $\operatorname{ad}_{[X,Y]}\operatorname{ad}_{[Z,W]}U=[[X,Y],[[Z,W],U]]=0,$
    as $\mathfrak{g}$ is $3$-step nilpotent.
    Furthermore, using that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, $\mathfrak{g}^{2}=[\mathfrak{g},\mathfrak{g}^{1}]\subset\mathfrak{z}$ and $\mathfrak{g}^{1},\mathfrak{g}^{1}]=0$,
    $2\nabla_{U}\operatorname{ad}_{[X,Y]}Z$ $=[U,[[X,Y],Z]]+[J_{1}U,[[X,Y],Z]]$
    $\ \ \ \ -J_{2}[U,J_{2}[[X,Y],Z]]+J_{3}[J_{1}U,J_{2}[[X,Y],Z]]=0,$
    and
    $2\operatorname{ad}_{[X,Y]}\nabla_{U}Z$ $=[[X,Y],[U,Z]]=0.$
    By Theorem 6.5, $\mathfrak{a}=\mathfrak{hol}(\nabla)$, concluding the proof. ∎
Theorem 6.1 and Proposition 7.1 highlight an unexpected rigidity in the holonomy algebra of the Obata connection. A related result is proved in [Ge], where the author shows that if the hypercomplex structure is abelian then the holonomy algebra of the Obata connection is abelian.
In the $3$-step nilpotent setting, as recalled above, all previously known examples were either Obata-flat or equipped with an abelian hypercomplex structure (the example constructed in this section is neither Obata-flat nor endowed with an abelian hypercomplex structure). Nevertheless, in every known case, the holonomy algebra of the Obata connection is abelian. This recurring behavior suggests that such rigidity might be a general feature of the nilpotent setting, motivating the following question:
Question 7.2.
Let $G$ be a nilpotent Lie group endowed with a left-invariant hypercomplex structure. Is it true that the holonomy algebra of the Obata connection is always abelian?