### v2

# Holonomy of the Obata connection on 2-step hypercomplex nilmanifolds

Abstract
We study the holonomy of the Obata connection on 2-step hypercomplex nilmanifolds. By explicitly computing the curvature tensor, we determine the conditions under which the Obata connection is flat, showing that this depends on the nilpotency step of each complex structure. In particular, we show that for 2-step hypercomplex nilmanifolds the holonomy algebra of the Obata connection is always an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$ and we prove that the $\mathbb{H}$-solvable conjecture holds in this case. Furthermore, we provide new examples of $k$-step nilpotent hypercomplex nilmanifolds, with arbitrary $k$, which are not Obata flat.
1 Introduction
Nilmanifolds, that is, compact quotients $\G\Gamma\backslash$ G of connected and simply connected nilpotent Lie groups $G$ by discrete subgroups $\Gamma$, have played a fundamental role in differential geometry and topology since the seminal works of Mal’cev and Mostow. They provide a natural class of spaces where algebraic and geometric properties interact tightly, offering a fair balance between algebraic manageability and rich geometric behaviour. In low dimensions, the latter allows classification results for various geometric structures (e.g., [Sal, FPS, Ug]), especially when these are left-invariant.
Invariant complex structures have been particularly investigated, largely due to the foundational work of Salamon [Sal], where it is shown that the existence of invariant complex structures $J$ on nilmanifolds admits special features: e.g., the existence of a so-called Salamon basis [Sal, Theorem 1.3], and the fact that the subspace $\mathfrak{g},\mathfrak{g}]+J[\mathfrak{g},\mathfrak{g}]$ is always a proper ideal of the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ [Sal, Corollary 1.4].
In the hypercomplex setting, many of these nice behaviours fail to carry over in a unified way. Each complex structure $J_{\alpha}$ ($\alpha=1,2,3$) may satisfy suitable compatibility conditions individually, but it is not clear whether these structures collectively induce analogous global properties. For instance, it is natural to ask whether the subspace
$\mathfrak{g}_{1}^{\mathbb{H}}:=[\mathfrak{g},\mathfrak{g}]+J_{1}[\mathfrak{g},\mathfrak{g}]+J_{2}[\mathfrak{g},\mathfrak{g}]+J_{3}[\mathfrak{g},\mathfrak{g}]$
is a proper ideal of $\mathfrak{g}$. This is an open question and has been explicitly conjectured in [Gor1, Gor2]. The difficulty in addressing this problem reflects the more rigid nature of hypercomplex geometry compared to the complex case.
A central geometric object in this setting is the Obata connection: the unique torsion-free connection that preserves all three complex structures. Its holonomy group $\operatorname{Hol}(\nabla)$ lies in $\operatorname{GL}(n,\mathbb{H})$ and plays a crucial role in understanding the geometry of hypercomplex manifolds [IP, SV]. In general, explicit descriptions of $\operatorname{Hol}(\nabla)$ are rare (see, for instance, [Sol, AGT, BFGV]). For nilmanifolds, however, it is known that the holonomy is contained in the commutator subgroup $\operatorname{SL}(n,\mathbb{H})$ [BDV, Corollary 3.3].
In this article, we investigate left-invariant hypercomplex structures on $2$-step nilpotent Lie groups. Surprisingly, all known examples in this class are Obata-flat.
This naturally leads to the question: Are all 2-step nilpotent hypercomplex Lie algebras Obata-flat? We answer this in the negative and clarify the role of the nilpotency step of each complex structure (see Sections 2 and 3 for details). In particular, we show that if the Obata connection is not flat, then each $J_{\alpha}$ must be 3-step nilpotent. No such examples had been previously known.
In Section 4, we construct examples of 2-step nilpotent hypercomplex Lie algebras in which one, two, or all three complex structures are 3-step nilpotent. Among these, Example 4.3 provides the first known example of a 2-step nilpotent hypercomplex Lie algebra with non-flat Obata connection. We also provide new examples of $k$-step nilpotent hypercomplex nilmanifolds, with arbitrary $k$, which are not Obata flat.
Turning to structural properties, we show the Gorgynian conjecture for $2$-step hypercomplex Lie algebras (Theorem 5.4).
In Section 6, we prove that the holonomy algebra of the Obata connection is always abelian for $2$-step nilpotent hypercomplex Lie groups, and we describe it explicitly (Theorem 6.1). This reveals a surprising rigidity: even when the Obata connection is non-flat, the holonomy remains severely restricted. We further show that the Obata holonomy group of a hypercomplex nilmanifold which is at most 2-step is trivial if and only if the nilmanifold is a torus (Theorem 6.9).
In the final part of the paper, we extend our analysis to the $3$-step nilpotent case, showing that, under suitable conditions, the holonomy of the Obata connection remains abelian (Proposition 7.2).
Acknowledgments. A. Andrada and M. L. Barberis were partially supported by CONICET, SECyT-UNC and ANPCyT (Argentina). B. Brienza was partially supported by Project PRIN 2022 “Real and complex manifolds: Geometry and Holomorphic Dynamics”, by GNSAGA (Indam) and by the INdAM - GNSAGA Project CUP E53C24001950001. The third author is grateful to A. Fino, G. Gentili, G. Gorginyan, and M. Verbitsky for many helpful discussions, and to the FAMAF for the warm hospitality during the preparation of this work. The authors would like to thank the anonymous referee for their detailed review of the manuscript.
2 Preliminaries
A complex structure on a differentiable manifold $M$ is an automorphism $J$ of the tangent bundle $TM$ satisfying $J^{2}=-I$ and the integrability condition $N_{J}(X,Y)=0$ for all vector fields $X,Y$ on $M$, where $N_{J}$ denotes the Nijenhuis tensor defined by
$N_{J}(X,Y)=[X,Y]+J([JX,Y]+[X,JY])-[JX,JY].$
The integrability of $J$ is equivalent to the existence of an atlas on $M$ whose transition functions are holomorphic maps [NN].
A hypercomplex structure on $M$ is given by a triple of complex structures $\{J_{\alpha}\}$, $\alpha=1,2,3$, satisfying the quaternionic relations
$J_{1}J_{2}=-J_{2}J_{1}=J_{3}.$ (2.1)
As a consequence, $M$ admits a 2-sphere family of complex structures $J_{y}=y_{1}J_{1}+y_{2}J_{2}+y_{3}J_{3},$ parametrized by points $y=(y_{1},y_{2},y_{3})$ in the unit sphere $S^{2}\subset\mathbb{R}^{3}$. It follows from (2.1) that at each point $p\in M$, the tangent space $T_{p}M$ carries a natural structure of a module over the quaternions $\mathbb{H}$. In particular, $\dim M\equiv 0\;\pmod{4}$.
Given a hypercomplex structure $\{J_{\alpha}\}$ on a manifold $M$, there exists a unique torsion-free connection $\nabla$ on $M$ that preserves each complex structure, i.e., satisfies $\nabla J_{\alpha}=0$ for $\alpha=1,2,3$. This connection, known as the Obata connection (see [Ob, Sol]), was shown in [Sol] to admit the following explicit expression:
$\nabla_{X}Y=\frac{1}{2}\bigl([X,Y]+J_{1}[J_{1}X,Y]-J_{2}[X,J_{2}Y]+J_{3}[J_{1}X,J_{2}Y]\bigr).$ (2.2)
More generally, by applying cyclic permutations $(\alpha,\beta,\gamma)$ of $(1,2,3)$, the Obata connection can also be written as:
$\nabla_{X}Y=\frac{1}{2}\bigl([X,Y]+J_{\alpha}[J_{\alpha}X,Y]-J_{\beta}[X,J_{\beta}Y]+J_{\gamma}[J_{\alpha}X,J_{\beta}Y]\bigr).$ (2.3)
It is known that $(M,\{J_{\alpha}\})$ is quaternionic (i.e., locally modelled over $\mathbb{H}^{n}$) if and only if the Obata connection $\nabla$ is flat [Ob, Theorem 11.2]. This highlights a fundamental difference with the complex case. For instance, the $K3$ surfaces and the homogeneous spaces $\operatorname{SU}(3)$ and $\operatorname{SO}(6)/\operatorname{SU}(2)$ provide examples of hypercomplex but not quaternionic manifolds [Boy, Joy].
Let $G$ be a Lie group with Lie algebra $\mathfrak{g}$. A hypercomplex structure $\{J_{\alpha}\}$ on $G$ is said to be left-invariant if left translations by elements of $G$ are hyperholomorphic, i.e., holomorphic with respect to each $J_{\alpha}$. In this case, the structure is completely determined by its value at the identity, which corresponds to a hypercomplex structure on $\mathfrak{g}$.
If $\Gamma$ is a lattice in $G$, any left-invariant hypercomplex structure on $G$ descends to a hypercomplex structure on the compact quotient $\G\Gamma\backslash$ G, which is then called invariant. In this setting, the natural projection $\GG\to\Gamma\backslash G$ is hyperholomorphic.
We recall the following definition.
Definition 2.1.
A Lie algebra $\mathfrak{g}$ is said to be nilpotent if its lower central series
$\mathfrak{g}=\mathfrak{g}^{0}\supset\mathfrak{g}^{1}\supset\cdots\supset\mathfrak{g}^{k}\supset\cdots,$
defined recursively by $\mathfrak{g}^{k}=[\mathfrak{g},\mathfrak{g}^{k-1}]$ for $k\geq 1$, eventually becomes zero; that is, there exists some $k\in\mathbb{N}$ such that
$\mathfrak{g}^{k}=0.$
The smallest such integer $k$ is called the nilpotency step of $\mathfrak{g}$.
Remark 2.2.
Equivalently, a Lie algebra $\mathfrak{g}$ is nilpotent if its upper central series
$0=\mathfrak{g}_{0}\subset\mathfrak{g}_{1}\subset\cdots\subset\mathfrak{g}_{k}\subset\cdots,$
defined recursively by
$\mathfrak{g}_{k}=\{X\in\mathfrak{g}\mid[X,\mathfrak{g}]\subset\mathfrak{g}_{k-1}\},\qquad k\geq 1,$
eventually coincides with $\mathfrak{g}$. Moreover, the smallest integer $k$ such that $\mathfrak{g}_{k}=\mathfrak{g}$ coincides with the nilpotency step defined via the lower central series.
Accordingly, a Lie algebra $\mathfrak{g}$ is said to be 2-step nilpotent if its commutator ideal $\mathfrak{g}^{1}=[\mathfrak{g},\mathfrak{g}]$ lies in the center $\mathfrak{z}$ of $\mathfrak{g}$. A Lie group $N$ is 2-step nilpotent if its Lie algebra $\mathfrak{g}$ is 2-step nilpotent.
Definition 2.3.
A 2-step nilmanifold is a compact quotient $\N\Gamma\backslash$ N, where $N$ is a connected, simply connected 2-step nilpotent Lie group and $\Gamma\subset N$ is a lattice.
It is well known that the existence of a lattice for a connected and simply connected nilpotent Lie group $N$ is equivalent to the existence of a rational basis of its Lie algebra $\mathfrak{g}$ [Mal], that is, a basis of $\mathfrak{g}$ for which the structure constants are rational numbers.
Let $(\mathfrak{g},J)$ be a nilpotent Lie algebra equipped with a complex structure $J$. In this context, it is natural to consider an ascending series adapted to $J$.
Definition 2.4 ([CFGU]).
The ascending series $\{\mathfrak{a}_{k}\}_{k\geq 0}$ (compatible with $J$) of $\mathfrak{g}$ is defined inductively by
$\mathfrak{a}_{0}=0,\quad\mathfrak{a}_{k}:=\{X\in\mathfrak{g}\ |\ [X,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\ \ \text{and}\ \ [JX,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\},\ k\geq 1.$ (2.4)
The complex structure $J$ is said to be nilpotent if there exists a $k\in\mathbb{N}$ such that $\mathfrak{a}_{k}=\mathfrak{g}$. The smallest such $k$ is called the nilpotency step of $J$.
If $\mathfrak{g}$ is $2$-step nilpotent, then any complex structure $J$ is nilpotent [Rol, Proposition 3.3], and the nilpotency step of $J$ is either $2$ or $3$ [GZZ, Theorem 1.3]. We recall from [B3] the following lemma.
Lemma 2.5 ([B3, Lemmas 3.3 and 4.1]).
Let $\mathfrak{g}$ be a $2$-step nilpotent Lie algebra endowed with a complex structure $J$. Then
  1. $(1)$
    $J$ is $2$-step nilpotent if and only if $J\mathfrak{g}^{1}\subset\mathfrak{z}$,
  2. $(2)$
    $J$ is $3$-step nilpotent if and only if there exists $X\in\mathfrak{g}^{1}$ such that $JX\notin\mathfrak{z}$.
  3. $(3)$
    $J\mathfrak{g}^{1}$ is an abelian ideal of $\mathfrak{g}$.
We include the proof of Statement (3) for further purposes. The proof is an application of the integrability of $J$. For any $X\in\mathfrak{g}$ and $Y\in\mathfrak{g}^{1}$, $N_{J}=0$ implies that
[JY,X]=J[JY,JX].[JY,X]=-J[JY,JX]. (2.5)
In particular, if $X\in J\mathfrak{g}^{1}$, then $[JY,X]=0$, as $\mathfrak{g}^{1}\subset\mathfrak{z}$. ∎
Remark 2.6.
Recall that a complex structure $J$ on a Lie algebra $\mathfrak{g}$ is called abelian if it satisfies $[JX,JY]=[X,Y]$ for all $X,Y\in\mathfrak{g}$ (see [BDM]). In the case when $\mathfrak{g}$ is 2-step nilpotent, an abelian complex structure is $2$-step nilpotent, since the center is $J$-invariant.
A hypercomplex structure $\{J_{\alpha}\}$ on a Lie algebra $\mathfrak{g}$ is called abelian if $J_{\alpha}$ is abelian for some $\alpha=1,2,3$, and it follows from [DF2, Lemma 3.1] that $J_{y}$ is abelian for each $y\in S^{2}$.
3 Computation of the Obata curvature
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. As already observed, the commutator ideal $\mathfrak{g}^{1}=[\mathfrak{g},\mathfrak{g}]$ lies in the center $\mathfrak{z}$ of $\mathfrak{g}$. This special feature allows for an explicit computation of the curvature tensor of the Obata connection.
Proposition 3.1.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. Then, for any $X,Y,Z\in\mathfrak{g}$ the curvature tensor of the Obata connection is given by
$4R(X,Y)Z=$ $-J_{\beta}[X,J_{\beta}[Y,Z]]+J_{\gamma}[J_{\alpha}X,J_{\beta}[Y,Z]]$
$-J_{\gamma}[X,J_{\beta}[J_{\alpha}Y,Z]]-J_{\beta}[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,Z]]$
$-[X,J_{\beta}[Y,J_{\beta}Z]]-J_{\alpha}[J_{\alpha}X,J_{\beta}[Y,J_{\beta}Z]]$
$+J_{\alpha}[X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]]-[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]]$
$+J_{\beta}[Y,J_{\beta}[X,Z]]-J_{\gamma}[J_{\alpha}Y,J_{\beta}[X,Z]]$
$+J_{\gamma}[Y,J_{\beta}[J_{\alpha}X,Z]]+J_{\beta}[J_{\alpha}Y,J_{\beta}[J_{\alpha}X,Z]]$ (3.1)
$+[Y,J_{\beta}[X,J_{\beta}Z]]+J_{\alpha}[J_{\alpha}Y,J_{\beta}[X,J_{\beta}Z]]$
$-J_{\alpha}[Y,J_{\beta}[J_{\alpha}X,J_{\beta}Z]]+[J_{\alpha}Y,J_{\beta}[J_{\alpha}X,J_{\beta}Z]]$
$-2[J_{\beta}[X,Y],J_{\beta}Z]+2J_{\gamma}[J_{\beta}[X,Y],J_{\alpha}Z],$
where $(\alpha,\beta,\gamma)$ is any cyclic permutation of $(1,2,3)$.
The proof is an application of the formula (2.3) and the integrability of $\{J_{\alpha}\}$. Using the formula (2.3) twice, the inclusion $\mathfrak{g}^{1}\subset\mathfrak{z}$ and that $\nabla_{X}J_{\alpha}Y=J_{\alpha}\nabla_{X}Y$, we compute:
$4\nabla_{X}\nabla_{Y}Z$ $=-J_{\beta}[X,J_{\beta}[Y,Z]]+J_{\gamma}[J_{\alpha}X,J_{\beta}[Y,Z]]$
$\quad-J_{\gamma}[X,J_{\beta}[J_{\alpha}Y,Z]]-J_{\beta}[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,Z]]$
$\quad-[X,J_{\beta}[Y,J_{\beta}Z]]-J_{\alpha}[J_{\alpha}X,J_{\beta}[Y,J_{\beta}Z]]$
$\quad+J_{\alpha}[X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]]-[J_{\alpha}X,J_{\beta}[J_{\alpha}Y,J_{\beta}Z]].$
Similarly, applying again (2.3) and using the inclusion $\mathfrak{g}^{1}\subset\mathfrak{z}$, we compute:
$2\nabla_{[X,Y]}Z$ $=J_{\alpha}[J_{\alpha}[X,Y],Z]+J_{\gamma}[J_{\alpha}[X,Y],J_{\beta}Z]$
$=[J_{\alpha}[X,Y],J_{\alpha}Z]+[J_{\beta}[X,Y],J_{\beta}Z]$
$\quad-[J_{\alpha}[X,Y],J_{\alpha}Z]-J_{\gamma}[J_{\beta}[X,Y],J_{\alpha}Z]$
$=[J_{\beta}[X,Y],J_{\beta}Z]-J_{\gamma}[J_{\beta}[X,Y],J_{\alpha}Z],$
where the second equality follows from the integrability of $J_{\alpha}$ on the first term and the integrability of $J_{\gamma}$ on the second term. ∎ As a consequence of Proposition 3.1 we obtain the following result. This is a generalization of [B1, Proposition 5.1], where it was shown that if $\mathfrak{z}$ is $J_{\alpha}$ invariant for all $\alpha$ then $\{J_{\alpha}\}$ is Obata flat.
Corollary 3.2.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. If any of the three complex structures $J_{1},J_{2},J_{3}$ is $2$-step nilpotent, then $(\mathfrak{g},\{J_{\alpha}\})$ is Obata flat.
If $J_{\beta}$ is $2$-step nilpotent then $J_{\beta}\mathfrak{g}^{1}\subset\mathfrak{z}$ (Lemma 2.5). The result then follows from Proposition 3.1, since each term contains an element of the form $\cdot,J_{\beta}\mathfrak{g}^{1}]$. ∎
Remark 3.3.
As a consequence of Corollary 3.2, we have that if the Obata connection associated to $(\mathfrak{g},\{J_{\alpha}\})$ is not flat, then $J_{\alpha}$ is $3$-step for all $\alpha$. Moreover, any complex structure $J_{y}$ in the hypercomplex 2-sphere is $3$-step.
4 Construction of non-flat 2-step hypercomplex nilmanifolds
To the best of our knowledge, in all known examples of $2$-step nilpotent hypercomplex Lie algebras $(\mathfrak{g},\{J_{\alpha}\})$, each complex structure $J_{\alpha}$ is $2$-step [B2, DF1, DF3, BD, LW, FG]. In particular, every known example of a $2$-step nilpotent Lie group equipped with a left-invariant hypercomplex structure is Obata flat (see Corollary 3.2).
According to [DF3, Section 3], in dimension $8$ any nilpotent Lie algebra admitting a hypercomplex structure is $2$-step nilpotent, and, moreover, each complex structure $J_{\alpha}$ is $2$-step.
In this section, we construct examples of $2$-step nilpotent hypercomplex Lie algebras for which one, two, or all of the complex structures are $3$-step. Moreover, we exhibit an example in which all three complex structures are $3$-step and the associated Obata connection is not flat.
The construction
Let $\mathfrak{n}$ be a $2$-step nilpotent Lie algebra and $\mathfrak{g}:=\mathfrak{n}\oplus\mathbb{R}^{r}$for some $r\in\mathbb{N}$.
Let $\mu\in\bigwedge^{2}\mathfrak{n}^{*}\otimes\mathbb{R}^{r}$ be a $2$-form on $\mathfrak{n}$ with values in $\mathbb{R}^{r}$ such that $\mathfrak{n}^{1}\subset\ker(\mu)$, where $\mathfrak{n}^{1}=[\mathfrak{n},\mathfrak{n}]$. We use $\mu$ to define a bracket $\cdot$$\cdot$$^{\prime}$ on $\mathfrak{g}$ as follows:
$[X,Y]^{\prime}=([X,Y],\mu(X,Y)),\ X,Y\in\mathfrak{n},$
$[X,Y]^{\prime}=0,\ X,Y\in\mathbb{R}^{r},$
$[X,Y]^{\prime}=0,\ X\in\mathfrak{n}\ \ \text{and}\ \ Y\in\mathbb{R}^{r}.$
Then, $(\mathfrak{g},[\cdot,\cdot]^{\prime})$ is a $2$-step nilpotent Lie algebra. We note that the condition $\mathfrak{n}^{1}\subset\ker(\mu)$ guarantees that the bracket $\cdot$$\cdot$$^{\prime}$ satisfies the Jacobi identity and that $\mathfrak{g}$ is $2$-step nilpotent.
Let us now assume that $\mathfrak{n}$ is a $2$-step nilpotent Lie algebra endowed with a hypercomplex structure $\{J_{\alpha}\}$, and let $(\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k},[\cdot,\cdot]^{\prime})$ be the $2$-step nilpotent Lie algebra constructed as above. Then $\mathfrak{g}$ carries a natural almost complex structure $J_{\alpha}^{\prime}=J_{\alpha}\oplus I_{\alpha}$, where $\{I_{\alpha}\}$ is the standard hypercomplex structure on $\mathbb{R}^{4k}$. The following result is a characterization of $N_{J^{\prime}_{\alpha}}=0$.
Lemma 4.1.
The almost complex structure $J^{\prime}_{\alpha}$ is integrable if and only if
$I_{\alpha}\mu(X,Y)=\mu(J_{\alpha}X,Y)+\mu(X,J_{\alpha}Y)+I_{\alpha}\mu(J_{\alpha}X,J_{\alpha}Y),\ X,Y\in\mathfrak{n}.$ (4.1)
Let us now assume that the $2$-step nilpotent hypercomplex Lie algebra $(\mathfrak{n},\{J_{\alpha}\})$ satisfies the condition that each $J_{\alpha}$ is $2$-step. If there exists $X\in\mathfrak{n}^{1}$ such that $\mu(J_{\alpha}X,\cdot)\neq 0$, then the corresponding complex structure $J^{\prime}_{\alpha}$ on $\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k}$ is $3$-step (Lemma 2.5). Since $X\in\mathfrak{n}^{1}$ we may write
$X=\sum_{j}\,[Y_{j},Z_{j}],\ \ \ Y_{j},Z_{j}\in\mathfrak{n}.$
We define
$X^{\prime}:=\sum_{j}\,[Y_{j},Z_{j}]^{\prime}=\sum_{j}([Y_{j},Z_{j}],\mu(Y_{j},Z_{j}))\in\mathfrak{g}^{1}.$
We claim that $J_{\alpha}^{\prime}X^{\prime}\notin\mathfrak{z}(\mathfrak{g})$. In fact, let $W\in\mathfrak{n}$ be such that $\mu(J_{\alpha}X,W)\neq 0$, then
$J^{\prime}_{\alpha}X^{\prime},W]^{\prime}=(0,\mu(J_{\alpha}X,W))\neq 0.$
We will use this procedure to construct new examples of $2$-step nilpotent hypercomplex Lie algebras where one, two or all three complex structures are $3$-step.
Examples
All $8$-dimensional hypercomplex nilpotent Lie algebras were classified in [DF3]. We consider the $8$-dimensional 22-step nilpotent Lie algebra $\mathfrak{n}_{8}$ defined by  [^1] We adopt the following notation: $e^{i,j}=e^{i}\wedge e^{j}$.
$de^{i}=0,\quad i=1,\dots,7,\qquad de^{8}=e^{1,2}-e^{3,4},$ (4.2)
which corresponds to the Lie algebra $\mathfrak{n}_{1}$ in [DF3, Section 3]. It was shown in [DF3, Section 3] (see also [DF1, Section 5]) that $\mathfrak{n}_{8}$ admits a hypercomplex structure $\{J_{\alpha}\}$, defined as follows:
$J_{1}e_{1}=e_{2},\ J_{1}e_{3}=e_{4},\ J_{1}e_{5}=e_{6},\ J_{1}e_{7}=e_{8},$
$J_{2}e_{1}=e_{3},\ J_{2}e_{2}=-e_{4},\ J_{2}e_{5}=e_{7},\ J_{2}e_{6}=-e_{8},$ (4.3)
$J_{3}e_{1}=e_{4},\ J_{3}e_{2}=e_{3},\ J_{3}e_{5}=e_{8},\ J_{3}e_{6}=e_{7}.$
In particular, since $\mathfrak{n}_{8}^{1}=\text{span}\{e_{8}\}$ and the center $\mathfrak{z}(\mathfrak{n}_{8})=\text{span}\{e_{5},e_{6},e_{7},e_{8}\}$, $J_{\alpha}$ is $2$-step for any $\alpha=1,2,3$.
Consider $\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4}$ and fix the standard basis $\{e_{9},e_{10},e_{11},e_{12}\}$ of $\mathbb{R}^{4}$. We point out that $\{I_{\alpha}\}$ are defined as
$I_{1}e_{9}=e_{10},\ I_{1}e_{11}=e_{12},$
$I_{2}e_{9}=e_{11},\ I_{2}e_{10}=-e_{12},$
$I_{3}e_{9}=e_{12},\ I_{3}e_{10}=e_{11}.$
Example 4.2.
We consider the following Lie bracket $\cdot$$\cdot$$^{\prime}$ on $\mathfrak{g}$ given by $\cdot,\cdot]^{\prime}=[\cdot,\cdot]+\mu$, where $\mu\in\bigwedge^{2}\mathfrak{n}_{8}^{*}\otimes\mathbb{R}^{4}$ is defined as
$\mu=e^{1,5}\otimes e_{9}+e^{2,5}\otimes e_{10}+e^{3,5}\otimes e_{11}+e^{4,5}\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. In particular, since $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{3}$ is $3$-step nilpotent, while $J^{\prime}_{1},J^{\prime}_{2}$ are $2$-step nilpotent.
To sum up, the $2$-step nilpotent Lie algebra $\mathfrak{g}$ defined by the structure equations
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{1,2}-e^{3,4},\\ &de^{9}=-e^{1,5},\ \ de^{10}=-e^{2,5},\\ &de^{11}=-e^{3,5},\ \ de^{12}=-e^{4,5},\end{split}$
can be endowed with a hypercomplex structure $\{J^{\prime}_{\alpha}\}$, where $J^{\prime}_{1},J^{\prime}_{2}$ are $2$-step nilpotent and $J^{\prime}_{3}$ is $3$-step nilpotent.
In order to construct a $2$-step nilpotent hypercomplex Lie algebra in which two of the complex structures are $3$-step nilpotent and one is $2$-step nilpotent, we consider the same underlying vector space $\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4}$ as above, but we modify the Lie bracket by changing the $2$-form $\mu$. Let us define on $\mathfrak{g}$ the Lie bracket $\cdot,\cdot]^{\prime}=[\cdot,\cdot]+\mu$, where $\mu\in\bigwedge^{2}\mathfrak{n}_{8}^{*}\otimes\mathbb{R}^{4}$ is now defined as
$\mu=\big(e^{1,5}+e^{1,6}\big)\otimes e_{9}+\big(e^{2,5}+e^{2,6}\big)\otimes e_{10}+\big(e^{3,5}+e^{3,6}\big)\otimes e_{11}+\big(e^{4,5}+e^{4,6}\big)\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. In particular, since $\mu(J_{2}e_{8},\cdot)=\mu(e_{6},\cdot)\neq 0$ and $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent, while $J^{\prime}_{1}$ is $2$-step nilpotent.
In summary, the $2$-step nilpotent Lie algebra $\mathfrak{g}$ defined by the structure equations
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ \ de^{8}=e^{1,2}-e^{3,4},\\ &de^{9}=-e^{1,5}-e^{1,6},\ \ de^{10}=-e^{2,5}-e^{2,6},\\ &de^{11}=-e^{3,5}-e^{3,6},\ \ de^{12}=-e^{4,5}-e^{4,6},\end{split}$
can be endowed with a hypercomplex structure $\{J^{\prime}_{\alpha}\}$, where $J^{\prime}_{1}$ is $2$-step nilpotent and $J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent.
We point out that both of examples constructed above are still Obata flat by Corollary 3.2.
Example 4.3.
In this last example we construct a $2$-step nilpotent hypercomplex Lie algebra where each complex structure is $3$-step. Let us consider again $\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4}$ with bracket $\cdot$$\cdot$$^{\prime}$ defined as $\cdot,\cdot]^{\prime}=[\cdot,\cdot]+\mu$ where $\mu\in\bigwedge^{2}\mathfrak{n}_{8}^{*}\otimes\mathbb{R}^{4}$ has the following definition:
$\mu=e^{5,6}\otimes e_{9}-e^{6,7}\otimes e_{11}+e^{5,7}\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. Since $\mu(J_{1}e_{8},\cdot)=\mu(-e_{7},\cdot)\neq 0$, $\mu(J_{2}e_{8},\cdot)=\mu(e_{6},\cdot)\neq 0$ and $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{1},J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent complex structures.
To sum up, the $2$-step nilpotent Lie algebra $\mathfrak{g}$ defined by the structure equations
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{1,2}-e^{3,4},\\ &de^{9}=-e^{5,6},\ \ de^{10}=0,\ \ \ de^{11}=e^{6,7},\ \ de^{12}=-e^{5,7},\end{split}$
admits a hypercomplex structure $\{J^{\prime}_{\alpha}\}$, where $J^{\prime}_{1},J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent. We claim that this hypercomplex Lie algebra is not Obata flat. Using Formula (2.2):
$\begin{split}R^{\prime}(e_{8},e_{1})e_{1}&=\nabla^{\prime}_{e_{8}}\nabla^{\prime}_{e_{1}}e_{1}=-\frac{1}{2}\nabla^{\prime}_{e_{8}}e_{7}=-\frac{1}{4}e_{9},\end{split}$
where we used that $\nabla^{\prime}_{e_{8}}e_{1}=0$ and $e_{1},e_{8}]=0$.
In Section 6, we will prove that if we consider the corresponding $2$-step nilpotent hypercomplex Lie group, then the holonomy algebra $\mathfrak{hol}(\nabla^{\prime})$ is abelian.
We point out that the $2$-step nilpotent Lie groups corresponding to the three Lie algebras constructed in this section admit lattices, since all the bases considered in the examples are rational.
Semidirect product construction
Next, we present a construction which is a slight generalization of [BF, Lemma 3.1].
Let $(\mathfrak{g},\{J_{\alpha}\})$ be an arbitrary hypercomplex Lie algebra and let $\{I_{\alpha}\}$ be the standard hypercomplex structure on $\mathbb{R}^{4k}$. Let $\rho:\mathfrak{g}\to\mathfrak{g}\mathfrak{l}(k,\mathbb{H})$ be a Lie algebra homomorphism, that is,
$\rho([X,Y])=[\rho(X),\rho(Y)],\quad\rho(X)I_{\alpha}={I_{\alpha}}_{\,}\rho(X),\;\;\text{ for all }X,\,Y\in\mathfrak{g},\;\alpha=1,2,3.$
Consider the Lie algebra $\mathfrak{h}:=\mathfrak{g}\ltimes_{\rho}\mathbb{R}^{4k}$, that is,
${[(X,V),(Y,W)]}=([X,Y],\rho(X)W-\rho(Y)V),\quad X,Y\in\mathfrak{g},\;V,W\in\mathbb{R}^{4k}.$
The Jacobi identity on $\mathfrak{h}$ follows from the fact that $\rho$ is a Lie algebra homomorphism. Note that $\{(0,V)\mid V\in\mathbb{R}^{4k}\}$ is an abelian ideal of $\mathfrak{h}$. Let $\widetilde{J}_{\alpha}$, $\alpha=1,2,3$, be the endomorphisms of $\mathfrak{h}$ defined by $\widetilde{J}_{\alpha}=J_{\alpha}\oplus I_{\alpha}$.
Proposition 4.4.
Let $(\mathfrak{g},\{J_{\alpha}\})$ and $(\mathfrak{h},\{\widetilde{J}_{\alpha}\})$ be as above, then $\{\widetilde{J}_{\alpha}\}$ is a hypercomplex structure on $\mathfrak{h}$. Moreover, the Obata connection $\widetilde{\nabla}$ of $\{\widetilde{J}_{\alpha}\}$ and the curvature tensor $\widetilde{R}$ of $\widetilde{\nabla}$ are given by:
$\widetilde{\nabla}_{(X,V)}(Y,W)$ $=$ $(\nabla_{X}Y,\rho(X)W),$
$\widetilde{R}\left((X,V),(Y,W)\right)(Z,U)$ $=$ $\left(R(X,Y)Z,0\right),\quad X,Y,Z\in\mathfrak{g},\;V,W,U\in\mathbb{R}^{4k},$
where $\nabla$ is the Obata connection of $\{J_{\alpha}\}$ and $R$ is the curvature tensor of $\nabla$. In particular, $\mathfrak{hol}(\widetilde{\nabla})\cong\mathfrak{hol}(\nabla)$.
Since $\{J_{\alpha}\}$ is integrable on $\mathfrak{g}$ and $\rho(X)$ commutes with $I_{\alpha}$ for all $X\in\mathfrak{g}$, $\alpha=1,2,3$, it turns out that $\{\widetilde{J}_{\alpha}\}$ is integrable on $\mathfrak{h}$.
The rest of the proof follows the lines of [BF, Lemma 3.1]. ∎
Lemma 4.5.
  1. $(1)$
    The center and commutator ideal of the Lie algebra $\mathfrak{h}$ are given by:
    $\mathfrak{z}(\mathfrak{h})$ $=$ $\left\{(X,V)\mid X\in\mathfrak{z}(\mathfrak{g})\cap\ker\rho,\;V\in\bigcap_{\begin{subarray}{c}Y\in\mathfrak{g}\end{subarray}}\ker\rho(Y)\right\},$
    $\mathfrak{h}^{1}$ $=$ $\left\{(X,V)\mid X\in\mathfrak{g}^{1},\;V\in\rho(\mathfrak{g})\mathbb{R}^{4k}\right\},$
    where $\rho(\mathfrak{g})\mathbb{R}^{4k}=\mathrm{span}\{\rho(X)V\mid X\in\mathfrak{g}\ \text{and}\ V\in\mathbb{R}^{4k}\}$.
  2. $(2)$
    If $\mathfrak{g}$ is $l$-step nilpotent and the set
    $\{j\in\mathbb{N}\mid\rho(X_{1})\cdots\rho(X_{j})=0,\ \text{for all}\ \ X_{1},\dots,X_{j}\in\mathfrak{g}\}$ (4.4)
    is non-empty, then $\mathfrak{h}$ is $r$-step nilpotent, with $r=\max\{l,m_{\rho}\}$, where $m_{\rho}$ is the minimum of the set defined in (4.4). In particular, if $\mathfrak{g}$ is $2$-step nilpotent and $\rho(X_{1})\cdot\rho(X_{2})=0$ for any $X_{1},X_{2}\in\mathfrak{g}$, then $\mathfrak{h}$ is $2$-step nilpotent.
  3. $(3)$
    If $\mathfrak{h}$ is 2-step nilpotent, then $\widetilde{J}_{\alpha}$ is 2-step nilpotent if and only if $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g})\cap\ker\rho$. In particular, if $J_{\alpha}$ is 3-step nilpotent, then $\widetilde{J}_{\alpha}$ is 3-step nilpotent.
The only nontrivial points are the second equality in (1) and statement (2).
The first claim follows from the computation
$\mathfrak{h}^{1}=[\mathfrak{h},\mathfrak{h}]=[(\mathfrak{g}\oplus\mathbb{R}^{4k},\,\mathfrak{g}\oplus\mathbb{R}^{4k})]=\mathfrak{g}^{1}\oplus\rho(\mathfrak{g})\,\mathbb{R}^{4k}.$
To prove the second statement, we recall that $\mathfrak{h}$ is $r$-step nilpotent if and only if $r$ is the smallest integer such that, for any $X_{i}\in\mathfrak{g}$ and $V_{i}\in\mathbb{R}^{4k}$,
$\prod_{i=1}^{r}\operatorname{ad}_{(X_{i},V_{i})}=0.$
We first observe that, for any $s\geq 1$,
$\left(\prod_{i=1}^{s}\operatorname{ad}_{(X_{i},0)}\right)(Y,W)=\left(\left(\prod_{i=1}^{s}\operatorname{ad}_{X_{i}}\right)Y,\;\left(\prod_{i=1}^{s}\rho(X_{i})\right)W\right).$
Since $\mathfrak{g}$ is $l$-step nilpotent and $m_{\rho}$ is minimal with the property above, this immediately implies that
$r\geq\mathrm{max}\{l,m_{\rho}\}.$
On the other hand, let $r=\max\{l,m_{\rho}\}$. Let $Y\in\mathfrak{g}$ and $W\in\mathbb{R}^{4k}$. Then, using the fact that
$\left(\prod_{i=1}^{r}\operatorname{ad}_{(X_{i},V_{i})}\right)(Y,W)\in\mathfrak{g}^{r}\oplus\rho(\mathfrak{g})^{\,r}(\mathbb{R}^{4k}),$
and the definition of $l$ and $m_{\rho}$, we conclude that
$\prod_{i=1}^{r}\operatorname{ad}_{(X_{i},V_{i})}=0.$
Therefore, $\mathfrak{h}$ is $r$-step nilpotent with $r=\max\{l,m_{\rho}\}$. ∎ New examples of Obata non flat 2-step nilpotent hypercomplex Lie algebras can be obtained by applying the semidirect product construction, as we show next.
Example 4.6.
Let $\mathfrak{g}$ be the Lie algebra from Example 4.3 with the hypercomplex structure considered there. We define $\rho:\mathfrak{g}\to\mathfrak{g}\mathfrak{l}(2,\mathbb{H})\subset\mathfrak{gl}(8,\mathbb{R})$ by:
$\rho(e_{1})=\begin{bmatrix}0_{4}&0_{4}\\ I_{4}&0_{4}\end{bmatrix},\qquad\quad\rho(e_{i})=0,\quad i\geq 2,$
[…]
Example 4.7.
Let $\mathfrak{g}$ be as in Example 4.3 and set $\rho:\mathfrak{g}\to\mathfrak{g}\mathfrak{l}(k,\mathbb{H})\subset\mathfrak{gl}(4k,\mathbb{R})$, $k\geq 3$, defined by:
$\rho(e_{1})=\left[\begin{array}[]{ccccc}0_{4}&0_{4}&\cdots&\cdots&0_{4}\\ I_{4}&0_{4}&0_{4}&\cdots&0_{4}\\ 0_{4}&I_{4}&0_{4}&\cdots&0_{4}\\ \vdots&\vdots&\ddots&\ddots&0_{4}\\ 0_{4}&0_{4}&0_{4}&I_{4}&0_{4}\end{array}\right]\in M_{4k}(\mathbb{R}),\qquad\rho(e_{i})=0,\quad i\geq 2.$
In this case, $\mathfrak{h}$ is $k$-step nilpotent and $\{\widetilde{J}_{\alpha}\}$ is not Obata flat. By Proposition 4.4, the holonomy algebra coincides with the holonomy algebra of the Example 4.3. In particular, it is abelian (see Section 6).
We note that the simply connected Lie groups corresponding to the Lie algebras from examples 4.6 and 4.7 admit lattices, since all the bases considered in the examples are rational.
5 $\mathbb{H}$-solvability of $2$-step nilpotent Lie algebras
Let $\mathfrak{g}$ be a Lie algebra endowed with a hypercomplex structure $\{J_{1},J_{2},J_{3}\}$. Given a vector subspace $W$ of $\mathfrak{g}$, we denote by $\mathbb{H}W$ the $\{J_{\alpha}\}$-invariant subspace $W+J_{1}W+J_{2}W+J_{3}W$.
Inspired by [Gor1, Definition 1.5], we introduce the following definition.
Definition 5.1.
Let $\mathfrak{g}$ be a Lie algebra endowed with a hypercomplex structure $\{J_{1},J_{2},J_{3}\}$. Define inductively $\{J_{\alpha}\}$-invariant subspaces:
$\mathfrak{g}_{0}^{\mathbb{H}}=\mathfrak{g},\qquad\mathfrak{g}_{k}^{\mathbb{H}}:=\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}},\quad k\geq 1.$
A hypercomplex Lie algebra $(\mathfrak{g},\{J_{\alpha}\})$ is called $\mathbb{H}$-solvable if the following sequence eventually becomes zero for some $k\in\mathbb{N}$:
$\mathfrak{g}_{1}^{\mathbb{H}}\supset\mathfrak{g}_{2}^{\mathbb{H}}\supset\cdots\supset\mathfrak{g}_{k-1}^{\mathbb{H}}\supset\mathfrak{g}_{k}^{\mathbb{H}}=0.$ (5.1)
The smallest such $k$ is called the $\mathbb{H}$-solvability step of $(\mathfrak{g},\{J_{\alpha}\})$.
Lemma 5.2.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be an arbitrary hypercomplex Lie algebra.
  1. $(1)$
    For any $k\in\mathbb{N}$, $\mathfrak{g}_{k}^{\mathbb{H}}$ is a subalgebra of $\mathfrak{g}$ and it is an ideal in $\mathfrak{g}_{k-1}^{\mathbb{H}}$.
  2. $(2)$
    If $(\mathfrak{g},\{J_{\alpha}\})$ is $\mathbb{H}$-solvable then $\mathfrak{g}$ is solvable.
We prove the first item. It suffices to show that $\mathfrak{g}_{k}^{\mathbb{H}}$ is an ideal of $\mathfrak{g}_{k-1}^{\mathbb{H}}$. Since $\mathfrak{g}^{1}\subset\mathfrak{g}_{1}^{\mathbb{H}}$, we have that $\mathfrak{g}_{1}^{\mathbb{H}}$ is clearly an ideal of $\mathfrak{g}_{0}^{\mathbb{H}}=\mathfrak{g}$. Assume by induction that $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is an ideal of $\mathfrak{g}_{k-2}^{\mathbb{H}}$. In particular, $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is a subalgebra of $\mathfrak{g}$. Then
$[\mathfrak{g}_{k}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}],\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset[\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset\mathfrak{g}_{k}^{\mathbb{H}}.$
To show the second, we work by induction by showing that $\mathfrak{g}_{k}^{\mathbb{H}}\supset D^{k}$, where $\{D^{k}\}_{k\geq 1}$ is the derived series of $\mathfrak{g}$.
For $k=1$, $\mathfrak{g}_{1}^{\mathbb{H}}\supset\mathfrak{g}^{1}=D^{1}$. Assume that $\mathfrak{g}_{k-1}^{\mathbb{H}}\supset D^{k-1}$. Then
$\mathfrak{g}_{k}^{\mathbb{H}}=\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\supset\mathbb{H}[D^{k-1},D^{k-1}]=\mathbb{H}D^{k}\supset D^{k}.$
Therefore, the $\mathbb{H}$-solvability of $(\mathfrak{g},\{J_{\alpha}\})$ implies the solvability of $\mathfrak{g}$. ∎ Let $(\mathfrak{g},J)$ be a nilpotent Lie algebra endowed with a complex structure $J$. It was shown by Salamon [Sal, Corollary 1.4] that the ideal $\mathfrak{g}^{1}+J\mathfrak{g}^{1}$ is a proper ideal of $\mathfrak{g}$. This naturally raises the question of whether an analogous property holds for nilpotent Lie algebras endowed with a hypercomplex structure.
The condition that $\mathfrak{g}_{1}^{\mathbb{H}}=\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1}$ is a proper ideal of $\mathfrak{g}$ is necessary to the $\mathbb{H}$-solvability of $(\mathfrak{g},\{J_{\alpha}\})$. This observation motivates the following conjecture:
Conjecture 5.3 ([Gor1]).
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a nilpotent hypercomplex Lie algebra. Then it is $\mathbb{H}$-solvable.
When the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ is $\mathbb{H}$-solvable, it turns out that, for a generic complex structure $J_{y}$ in the hypercomplex $2$-sphere, the complex nilmanifold $(\Gamma\backslash G,J_{y})$ does not admit any complex curves [Gor1].
The conjecture holds for Lie algebras of hypercomplex nilmanifolds with flat Obata connection [Gor2], as well as for Lie algebras of abelian hypercomplex nilmanifolds [AV]. The filtration (5.1) corresponds to an iterated hypercomplex toric bundle [Gor1], and in the latter case it was shown in [AV, Proposition 4.5] that such iterated torus bundle terminates at a point. Moreover, for a general complex structure $J_{y}$, the associated complex nilmanifold $(\Gamma\backslash G,J_{y})$ not only admits no complex curves (as a consequence of [Gor1]), but in fact all of its subvarieties are trianalytic.
Here we prove the conjecture for $2$-step nilpotent Lie algebras.
Theorem 5.4.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a $2$-step nilpotent hypercomplex Lie algebra. Then $(\mathfrak{g},\{J_{\alpha}\})$ is $\mathbb{H}$-solvable. In particular, $\mathfrak{g}_{1}^{\mathbb{H}}$ is a proper ideal of $\mathfrak{g}$ and the $\mathbb{H}$-solvability step is at most $3$.
Let us consider
$\mathfrak{g}_{2}^{\mathbb{H}}=\mathbb{H}[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}].$
Using that $\mathfrak{g}^{1}\subset\mathfrak{z}$, and Lemma 2.5 (3), we obtain that $J_{\alpha}\mathfrak{g}^{1},J_{\alpha}\mathfrak{g}^{1}]=0$, and
$[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}]=$ $[\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1},\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1}]$
$=$ $[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}].$
Therefore,
$\mathbb{H}[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}]=$ $\ [J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]$
$+J_{1}([J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}])$
$+J_{2}([J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}])$
$+J_{3}([J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]+[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]).$
Since by Equation (2.5) we have that $J_{\alpha}[J_{\alpha}\mathfrak{g}^{1},J_{\beta}\mathfrak{g}^{1}]\subset\mathfrak{g}^{1}$, we may write:
$\mathfrak{g}^{\mathbb{H}}_{2}=\mathbb{H}[\mathfrak{g}_{1}^{\mathbb{H}},\mathfrak{g}_{1}^{\mathbb{H}}]$ $\subset\mathfrak{g}^{1}+J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]+J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}].$
We claim that $J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]$. Using the quaternionic identities and Equation (2.5) with $J_{1}$, we have:
$J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}J_{1}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}],$
and, similarly, using the integrability of $J_{2}$,
$J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]=J_{1}J_{2}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]=J_{1}[J_{3}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}].$
Thus, all three summands coincide and we obtain that $\mathfrak{g}^{\mathbb{H}}_{2}\subset\mathfrak{g}^{1}+J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]\subset\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}$. Therefore,
$\mathfrak{g}_{3}^{\mathbb{H}}=\mathbb{H}[\mathfrak{g}^{\mathbb{H}}_{2},\mathfrak{g}^{\mathbb{H}}_{2}]\subset\mathbb{H}[J_{1}\mathfrak{g}^{1},J_{1}\mathfrak{g}^{1}]=0,$
since $J_{1}\mathfrak{g}^{1}$ is abelian by Lemma 2.5. ∎
6 Holonomy of 2-step hypercomplex nilmanifolds
We prove a structural result about the holonomy algebra of the Obata connection on $2$-step nilpotent Lie groups endowed with invariant hypercomplex structures.
Theorem 6.1.
Let $G^{4n}$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. Then the holonomy Lie algebra $\mathfrak{hol}(\nabla)$ is abelian, where $\nabla$ is the Obata connection of $(G,\{J_{\alpha}\})$.
Moreover, $A\cdot B=0$ for any $A,B\in\mathfrak{hol}(\nabla)$, therefore, $\mathfrak{hol}(\nabla)$ is a commutative associative algebra with trivial product.
The proof is divided into $4$ steps. We start by proving
  • Step 1: for any $X,Y,Z\in\mathfrak{g}$, $R(X,Y)Z\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$.
    For $(\alpha,\beta,\gamma)=(1,2,3)$, each term appearing in Formula (3.1) falls into one of four distinct types:
    1. 1.
      $J_{2}[\cdot,J_{2}[\cdot,\cdot]]=[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, where the equality follows from Equation (2.5),
    2. 2.
      $\cdot,J_{2}[\cdot,\cdot]]\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, as $J_{2}\mathfrak{g}^{1}$ is an ideal by Lemma 2.5,
    3. 3.
      $J_{3}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows from Equation (2.5),
    4. 4.
      $J_{1}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows again from Equation (2.5).
    We set $\mathfrak{h}:=\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. It is not difficult to note that $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant and $\mathfrak{h}\subset\mathfrak{g}_{1}^{\mathbb{H}}\subsetneq\mathfrak{g}$.
  • Step 2: for any $X\in\mathfrak{h}$, $\nabla X=0$.
    We use the following:
    Lemma 6.2.
    For any $X\in\mathfrak{h}$ and $Y\in\mathfrak{g}$,
    $Y,X]+J_{1}[J_{1}Y,X]=0$ (6.1)
    • Proof.
      Let $X\in\mathfrak{h}$. Then $X=X_{1}+X_{2}$, where $X_{1}\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$ and $X_{2}\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. Since $X_{1}\in\mathfrak{g}^{1}$,
      $[Y,X]+J_{1}[J_{1}Y,X]$ $=[Y,X_{2}]+J_{1}[J_{1}Y,X_{2}]$
      $=[Y,X_{2}]-[Y,X_{2}]+[J_{1}Y,J_{1}X_{2}]-J_{1}[Y,J_{1}X_{2}]$
      $=0$
      since $J_{1}X_{2}\in\mathfrak{g}^{1}\subset\mathfrak{z}$, where we have used the integrability of $J_{1}$ in the second equality. ∎
    Let $Y\in\mathfrak{g}$ and $X\in\mathfrak{h}$. Applying Formula (2.2)
    $2\nabla_{Y}X$ $=[Y,X]+J_{1}[J_{1}Y,X]-J_{2}[Y,J_{2}X]+J_{3}[J_{1}Y,J_{2}X]$
    $=[Y,X]+J_{1}[J_{1}Y,X]-J_{2}([Y,J_{2}X]+J_{1}[J_{1}Y,J_{2}X])$
    $=0,$
    where we used Lemma 6.2 twice and the $\{J_{\alpha}\}$-invariance of $\mathfrak{h}$.
  • Step 3: Expression for the curvature endomorphisms. Using Steps 1 and 2, we conclude that for any $X,Y\in\mathfrak{g}$,
    $R(X,Y)\in\bigl\{A\in\mathfrak{gl}(n,\mathbb{H})\mid\operatorname{Im}(A)\subset\mathfrak{h}\subset\ker(A)\bigr\}.$
    Moreover, since $R(X,Y)^{2}=0$, we have that $R(X,Y)\in\mathfrak{sl}(n,\mathbb{H})$ and therefore
    $R(X,Y)\in\bigl\{A\in\mathfrak{sl}(n,\mathbb{H})\mid\operatorname{Im}(A)\subset\mathfrak{h}\subset\ker(A)\bigr\}=:\mathfrak{a}.$
  • Step 4: Application of the Ambrose-Singer Theorem.
    We recall the following version of the Ambrose-Singer Theorem for left-invariant connections.
    Theorem 6.3.
    [Al] Let $\nabla$ be a left-invariant affine connection on the Lie group $G$, and let $\mathfrak{g}$ denote the Lie algebra of $G$. Then the holonomy algebra $\mathfrak{hol}(\nabla)$ based at the identity element $e\in G$ is the smallest subalgebra of $\mathfrak{gl}(\mathfrak{g})$ containing the curvature endomorphisms $R(X,Y)$ for any $X,Y\in\mathfrak{g}$, and closed under commutators with the left multiplication operators $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$.
    Let us consider $\text{span}\{R(X,Y)\mid X,Y\in\mathfrak{g}\}$. As already observed in Step $3$,
    $\text{span}\{R(X,Y)\mid X,Y\in\mathfrak{g}\}\subset\mathfrak{a}.$
    We point out that $\mathfrak{a}$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$.
    Let $A\in\mathfrak{a}$ and let $X\in\mathfrak{g}$.
    $A,\nabla_{X}]\in\mathfrak{a}.$
    In fact, let $Y\in\mathfrak{g}$ and $Y^{\prime}\in\mathfrak{h}$:
    $A\nabla_{X}Y-\nabla_{X}AY\subset A(\mathfrak{g}_{1}^{\mathbb{H}})-\nabla_{X}\mathfrak{h}=A(\mathfrak{g}_{1}^{\mathbb{H}})\subset\mathfrak{h},$
    that is, $\mathrm{Im}([A,\nabla_{X}])\subset\mathfrak{h}$, and
    $A\nabla_{X}Y^{\prime}-\nabla_{X}AY^{\prime}=0,$
    implying that $\mathfrak{h}\subset\ker([A,\nabla_{X}])$. Moreover, $A,\nabla_{X}]\in\mathfrak{sl}(n,\mathbb{H})$ since $\mathfrak{sl}(n,\mathbb{H})$ is an ideal of $\mathfrak{gl}(n,\mathbb{H})$.
    Since $\mathfrak{a}$ is closed under commutators with $\nabla_{\mathfrak{g}}$, by Theorem 6.3 we get that $\mathfrak{hol}(\nabla)\subset\mathfrak{a}$. In particular, $\mathfrak{a}$ is abelian and has trivial product.
Example 6.4.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be the $2$-step nilpotent hypercomplex Lie algebra constructed in Example 4.3 and let us consider the corresponding $2$-step nilpotent hypercomplex Lie group $G$. As already observed, the Obata connection $\nabla$ on $G$ is not flat. A direct computation shows that the holonomy algebra $\mathfrak{hol}(\nabla)$ is given by
$\mathfrak{hol}(\nabla)=\left\{\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr q&p\cdot j&0\end{array}\end{pmatrix}\in\mathfrak{sl}(3,\mathbb{H})\ |\ q\in\mathbb{H},\ p\in\mathbb{R}\right\},$
where we are considering $\mathfrak{g}=\text{span}_{\mathbb{H}}\{e_{1},e_{5},e_{9}\}$.
This provides an explicit example of a holonomy algebra arising from a non-flat Obata connection on a $2$-step nilmanifold such that $\mathfrak{hol}(\nabla)$ is abelian and non-trivial.
The next lemma is one of the key ingredients in the proof of Theorem 6.9 below.
Lemma 6.5.
Let $G^{4n}$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. Then for any $X,Y\in\mathfrak{g}$, $\nabla_{X}Y\in\mathfrak{g}_{1}^{\mathbb{H}}$ and $\nabla_{X}\mathfrak{g}_{1}^{\mathbb{H}}\subset\mathfrak{h}$. In particular, for any $\in\mathfrak{g}, X\nabla_{X}$ is nilpotent.
The first claim is straightforward by Formula (2.2).
For the second claim, since $\nabla_{X}J_{\alpha}=J_{\alpha}\nabla_{X}$ and $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant, it suffices to prove that $\nabla_{X}\mathfrak{g}^{1}\subset\mathfrak{h}$. Applying Formula (2.2), we see that for any $Y\in\mathfrak{g}^{1}$:
$\nabla_{X}Y=\frac{1}{2}\big(-J_{2}[X,J_{2}Y]+J_{3}[J_{1}X,J_{2}Y]\big).$
With the same argument used in the proof of Step 1 in Theorem 6.1, the first term lies in $\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, while the second term lies in $J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. The proof follows by the definition of $\mathfrak{h}$.
The fact that $\nabla_{X}$ is nilpotent follows since $(\nabla_{X})^{3}=0.$ In fact for any $Y\in\mathfrak{g}$
$\nabla_{X}(\nabla_{X}(\nabla_{X}Y))\in\nabla_{X}(\nabla_{X}\mathfrak{g}_{1}^{\mathbb{H}})\subset\nabla_{X}\mathfrak{h}=0,$
by Step 2 in the proof of Theorem 6.1. ∎
6.1 Trivial Obata holonomy
Let $G$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. If the Obata connection is flat, then the full holonomy group is discrete. A natural question is when the holonomy group is trivial; in this section we prove in Theorem 6.9 below that it is never the case.
In order to prove this, we recall some useful facts on affine connections. Let $M$ be a connected $n$-dimensional manifold with an affine connection $\nabla$. We define the following subspace of the space $\mathfrak{X}(M)$ of smooth vector fields on $M$:
$\mathcal{P}^{\nabla}=\{X\in\mathfrak{X}(M)\mid\nabla X=0\}.$ (6.2)
The subspace $\mathcal{P}^{\nabla}$ is in general neither finite dimensional as an $\mathbb{R}$-vector space nor closed under the Lie bracket of vector fields. However, under additional assumptions both conditions are satisfied, as the following well known lemma shows (see [Wo, Proposition 2.2] for the first item and [Wo, p. 323] for the second one).
Lemma 6.6 ([Wo]).
Let $M$ be a connected $n$-dimensional manifold with an affine connection $\nabla$.
  1. $(1)$
    If the holonomy group $\operatorname{Hol}(\nabla)$ is trivial, then $\mathcal{P}^{\nabla}$ is an $n$-dimensional real vector space.
  2. $(2)$
    The space $\mathcal{P}^{\nabla}$ is a Lie subalgebra of $\mathfrak{X}(M)$ if and only if the torsion $T$ of $\nabla$ is parallel.
In case the manifold admits an almost complex structure $J$ and an affine connection $\nabla$ with trivial holonomy, we will make use of the following result.
Lemma 6.7 ([ABD, Lemma 3.2]).
Let $M$ be a connected manifold with an almost complex structure $J$. Assume that there exists an affine connection $\nabla$ on M such that $\operatorname{Hol}(\nabla)$ is trivial. Then the following conditions are equivalent:
  1. $(1)$
    $\nabla J=0$;
  2. $(2)$
    the space $\mathcal{P}^{\nabla}$ of parallel vector fields on $M$ is $J$-invariant, that is, $J(\mathcal{P}^{\nabla})=\mathcal{P}^{\nabla}$.
Any Lie group $H$ admits a natural left-invariant connection with trivial holonomy: it is defined by $\nabla^{0}_{X}Y=0$ for all $X,Y$ left-invariant vector fields on $H$, and it is known as the ()(-)-connection (see, for instance, [KN, Chapter X, Proposition 2.12]). Note that its torsion is given by $T^{0}(X,Y)=-[X,Y]$ for $X,Y$ left-invariant vector fields on $H$. Moreover, $\nabla^{0}$ is flat, complete, and clearly $\nabla^{0}T^{0}=0$ by definition of $\nabla^{0}$.
If $\Lambda\subset H$ is any discrete subgroup of $H$, then $\nabla^{0}$ on $H$ induces a unique connection, still denoted $\nabla^{0}$, on $\H\Lambda\backslash$ H such that the projection $\H\pi:H\to\Lambda\backslash$ H is an affine local diffeomorphism. This connection $\nabla^{0}$ on $\H\Lambda\backslash$ H is geodesically complete, has trivial holonomy and its torsion is parallel.
If $J_{0}$ is a left-invariant complex structure on $H$ then, by definition of $\nabla^{0}$, it follows that $\nabla^{0}J_{0}=0$. The complex structure $J_{0}$ on $H$ induces a unique complex structure, still denoted $J_{0}$, on $\H\Lambda\backslash$ H such that $\pi$ is a local biholomorphism. Since $\pi$ is also affine, we have that $\nabla^{0}J_{0}=0$ on $\H\Lambda\backslash$ H. Therefore, $(\Lambda\backslash H,J_{0})$ carries a geodesically complete connection $\nabla^{0}$ with trivial holonomy and parallel torsion such that $\nabla^{0}J_{0}=0$. See [ABD] for further details.
Theorem 6.8 ([ABD, Theorem 4.1]).
The triple $(M,J,\nabla)$ where $M$ is a connected manifold endowed with a complex structure $J$ and a connection $\nabla$ with $\nabla J=0$ and trivial holonomy is affinely biholomorphic to a triple $(\Lambda\backslash H,J_{0},\nabla^{0})$ as above if and only if $\nabla$ is geodesically complete and its torsion is parallel.
The Lie group $H$ is the connected and simply connected Lie group with Lie algebra $\mathcal{P}^{\nabla}$ defined in (6.2), and $J_{0}$ is the complex structure on $\H\Lambda\backslash$ H induced by the restriction of $J$ to $\mathcal{P}^{\nabla}$.
As a consequence of Lemmas 6.5, 6.6 and 6.7 and Theorem 6.8 we obtain the main result of this section.
Theorem 6.9.
Let $(\Gamma\backslash G,\{J_{\alpha}\})$ be an at most 2-step nilmanifold endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$ such that the Obata connection $\nabla$ is flat. Then the holonomy group $\operatorname{Hol}(\nabla)$ is trivial if and only if $\G\Gamma\backslash$ G is a torus.
In particular, if $(\Gamma\backslash G,\{J_{\alpha}\})$ is a $2$-step nilmanifold endowed with an invariant hypercomplex structure such that the Obata connection $\nabla$ is flat, then the holonomy group $\operatorname{Hol}(\nabla)$ is never trivial.
The implication from right to left is clear. Let us now prove the converse.
The Obata connection $\nabla$ satisfies $\nabla J_{\alpha}=0$ for all $\alpha$ and has vanishing (and hence parallel) torsion. By assumption, $\operatorname{Hol}(\nabla)$ is trivial, therefore, Lemma 6.6 implies that the space $\mathcal{P}^{\nabla}$ of smooth parallel vector fields on $\G\Gamma\backslash$ G is a $4n$-dimensional Lie algebra, with $4n:=\dim\Gamma\backslash G=\dim G$, and it follows from Lemma 6.7 that $J_{\alpha}(\mathcal{P}^{\nabla})=\mathcal{P}^{\nabla}$ for all $\alpha$. We denote by $H$ the connected and simply connected Lie group with Lie algebra $\mathcal{P}^{\nabla}$.
On the other hand, it follows from Lemma 6.5 that, for any $X\in\mathfrak{g}$, the endomorphism $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$ is nilpotent. As the Obata connection $\nabla$ is flat and torsion-free, it follows from [Kim, Theorem 2.2 and Proposition 1.2] that $\nabla$ is geodesically complete on $G$, hence on $\G\Gamma\backslash$ G.
Since $\nabla$ is flat, torsion-free, geodesically complete and $\nabla J_{\alpha}=0$ for all $\alpha$, we may apply Theorem 6.8 to conclude that $(\Gamma\backslash G,J_{\alpha},\nabla)$ is affinely biholomorphic to $(\Lambda\backslash H,J_{\alpha}^{0},\nabla^{0})$ for each $\alpha$, where $\Lambda$ is a lattice in $H$, $J_{\alpha}^{0}$ is the complex structure induced by $J_{\alpha}$ on $\H\Lambda\backslash$ H and $\nabla^{0}$ is the ()(-)-connection on $\H\Lambda\backslash$ H.
As $\nabla$ is torsion free, so is $\nabla^{0}$, therefore, we have that $[X,Y]=0$ for any $X,Y$ vector fields on $\H\Lambda\backslash$ H induced by left-invariant vector fields on $H$. In other words, $[X,Y]=0$ for any $X,Y\in\mathcal{P}^{\nabla}$. In particular, $H$ is abelian and therefore $\H\Lambda\backslash$ H is a torus.
Since $\G\Gamma\backslash$ G is diffeomorphic to $\H\Lambda\backslash$ H and both $G$ and $H$ are nilpotent, then $G$ is isomorphic to $H$ as a Lie group [Mal]. It then follows that $G$ is abelian, and so $\G\Gamma\backslash$ G is also a torus. ∎
Remark 6.10.
Theorem 6.9 still holds true if instead of considering a 2-step nilmanifold we assume that $\G\Gamma\backslash$ G is a solvmanifold with $G$ a completely solvable almost abelian Lie group. Indeed, in this case the Obata connection is automatically flat, according to [AB1, Proposition 3.7], and it is complete, due to [AB1, Corollary 3.8]. For the last paragraph, we use the Saito rigidity theorem [Sai], which extends the result of Mal’cev from nilpotent to completely solvable Lie groups.
7 Finale
If $(\mathfrak{g},\{J_{\alpha}\})$ is a hypercomplex Lie algebra such that $\mathfrak{g}$ is $2$-step nilpotent, we have seen that the Obata connection is flat when at least one complex structure $J_{\alpha}$ satisfies
$J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}.$ (7.1)
We now investigate to what extent this condition can be extended to Lie algebras of higher nilpotency step. It turns out that (7.1) forces the Lie algebra to be at most $3$-step nilpotent.
Lemma 7.1.
  1. $(1)$
    Let $(\mathfrak{g},J)$ be a Lie algebra endowed with a complex structure $J$ satisfying condition (7.1). Then $\mathfrak{g}$ is at most $3$-step nilpotent and $\mathfrak{g}^{1}$ is abelian.
  2. $(2)$
    Let $(\mathfrak{g},\{J_{1},J_{2},J_{3}\})$ be a hypercomplex Lie algebra and assume that $(\alpha,\beta,\gamma)$ is a cyclic permutation of $(1,2,3)$ such that $J_{\alpha}$ and $J_{\beta}$ satisfy (7.1). Then $J_{\gamma}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g}^{1})$ and $\nabla_{X}Y=0$ for any $Y\in\mathfrak{g}^{1}$, where $\nabla$ is the Obata connection.
In case (1), by the integrability of $J$ and the assumption $J\mathfrak{g}^{1}\subset\mathfrak{z}$, one obtains
$X,Y],Z]=-J[[X,Y],JZ],\quad\text{for any }X,Y,Z\in\mathfrak{g},$ (7.2)
which shows that $\mathfrak{g}^{2}\subset\mathfrak{z}$, and hence $\mathfrak{g}^{3}=[\mathfrak{g}^{2},\mathfrak{g}]=0$. Note that by Equation (7.2), $\mathfrak{g}^{1},\mathfrak{g}^{1}]=0$.
To prove (2), we may assume that $J_{1}$ and $J_{2}$ satisfy (7.1). We start by showing that $J_{3}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g}^{1})$, that is, $X,Y],J_{3}[U,Z]]=0$ for any $X,Y,U,Z\in\mathfrak{g}$. Indeed, using the integrability of $J_{1}$:
$[[X,Y],J_{3}[U,Z]]$ $=-J_{1}[J_{1}[X,Y],J_{3}[U,Z]]-J_{1}[[X,Y],J_{1}J_{3}[U,Z]]+[J_{1}[X,Y],J_{1}J_{3}[U,Z]]$
$=-J_{1}[J_{1}[X,Y],J_{3}[U,Z]]+J_{1}[[X,Y],J_{2}[U,Z]]-[J_{1}[X,Y],J_{2}[U,Z]]$
$=0,$
since $J_{1}\mathfrak{g}^{1}\subset\mathfrak{z}$ and $J_{2}\mathfrak{g}^{1}\subset\mathfrak{z}$. The second part of (2) follows by computing $\nabla_{X}[V,W]$ for any $X,V,W\in\mathfrak{g}$ using (2.2) and by applying the fact that both $J_{1}$ and $J_{2}$ satisfy (7.1) and (7.2). ∎
Since in the $2$-step case the corresponding Obata connection is flat (see Corollary 3.2), we will focus our attention on the $3$-step case. We observe that the hypercomplex structure must be non-abelian. Indeed, if $\{J_{\alpha}\}$ is abelian and $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for at least one $\alpha$, then
$\mathfrak{g}^{1}\subset J_{\alpha}\mathfrak{z}=\mathfrak{z},$
which forces $\mathfrak{g}$ to be $2$-step nilpotent.
Examples of 3-step nilpotent Lie algebras with hypercomplex structures have been given in [DF2, BD, AB2]. Among these, the only examples with non-abelian hypercomplex structures are those in [AB2]; however, those are Obata flat and none of them satisfies the hypothesis of Lemma 7.1 (2).
We proved in Theorem 6.1 that for hypercomplex 2-step nilmanifolds the holonomy of the Obata connection is an abelian Lie algebra, which is also a commutative associative algebra with trivial product. This theorem requires no additional assumption on the hypercomplex structure. Next, we prove a similar result without assuming 2-step nilpotency, but requiring an extra condition on the hypercomplex structure, namely, that $\{J_{1},J_{2},J_{3}\}$ satisfies the hypothesis of Lemma 7.1 (2).
Proposition 7.2.
Let $(\mathfrak{g},\{J_{1},J_{2},J_{3}\})$ be a $4n$-dimensional hypercomplex Lie algebra and assume that $(\alpha,\beta,\gamma)$ is a cyclic permutation of $(1,2,3)$ such that $J_{\alpha}$ and $J_{\beta}$ satisfy (7.1). Then:
  1. $(1)$
    the curvature tensor of the Obata connection is given by $R(X,Y)=-\operatorname{ad}_{[X,Y]}$, hence $(\mathfrak{g},\{J_{\alpha}\})$ is non-flat if and only if $\mathfrak{g}$ is $3$-step nilpotent;
  2. $(2)$
    the holonomy algebra $\mathfrak{hol}(\nabla)$ of the Obata connection $\nabla$ on $G$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$, where $G$ is the connected and simply connected Lie group with Lie algebra $\mathfrak{g}$. Moreover, $A\cdot B=0$ for any $A,B\in\mathfrak{hol}(\nabla)$, therefore, $\mathfrak{hol}(\nabla)$ is a commutative associative algebra with trivial product.
To prove (1), we first show that $\nabla_{X}\nabla_{Y}Z=0$. We start by computing $\nabla_{Y}Z$ applying Formula (2.2). Since $\nabla_{X}$ commutes with $J_{\alpha}$ for all $\alpha$, Lemma 7.1 (2) implies that $\nabla_{X}\nabla_{Y}Z=0$. Therefore,
$R(X,Y)Z=-\nabla_{[X,Y]}Z=-\nabla_{Z}[X,Y]+[Z,[X,Y]]=-[[X,Y],Z],$
where the second equality follows from the fact that $\nabla$ is torsion-free.
In order to prove (2), set
$\mathfrak{b}:=\text{span}\{\operatorname{ad}_{[X,Y]}\mid X,Y\in\mathfrak{g}\}.$
Next, we show that $\mathfrak{b}$ is a commutative associative algebra with trivial product and that $\mathfrak{hol}(\nabla)=\mathfrak{b}$. In fact,
$\operatorname{ad}_{[X,Y]}\circ\operatorname{ad}_{[Z,W]}U=[[X,Y],[[Z,W],U]]=0,$
as $\mathfrak{g}^{1}$ is abelian (Lemma 7.1 (1)). Therefore, $\mathfrak{b}$ is a commutative associative algebra with trivial product.
Next, we show that $\mathfrak{hol}(\nabla)=\mathfrak{b}$. From Lemma 7.1 (2), we have that
$\nabla_{Z}\operatorname{ad}_{[X,Y]}=0.$
On the other hand,
$2\operatorname{ad}_{[X,Y]}\nabla_{Z}U$ $=\operatorname{ad}_{[X,Y]}J_{3}[{J_{1}Z},J_{2}U]=0,$
where the last equality follows from $J_{3}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g}^{1})$ (Lemma 7.1 (2)). By Theorem 6.3, $\mathfrak{hol}(\nabla)=\mathfrak{b}$, which implies that $\mathfrak{hol}(\nabla)$ is an abelian Lie subalgebra of $\mathfrak{sl}(n,\mathbb{H})$, which is also an associative commutative algebra with trivial product, concluding the proof. ∎
With techniques similar to those in Section 4, we can construct an example of a $3$-step nilpotent Lie algebra endowed with a hypercomplex structure $\{J_{\alpha}\}$ such that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for each $\alpha$.
Example 7.3.
Let $\mathfrak{g}$ be the $3$-step nilpotent Lie algebra defined by the structure equations
$\begin{split}&de^{i}=0,\ i=1,\dots,4,\ \ de^{5}=e^{1,2}-e^{3,4},\\ &de^{i}=0,\ i=6,7,8,\ \ de^{9}=e^{1,3}-e^{4,2},\\ &de^{i}=0,\ i=10,11,12,\ \ de^{13}=-e^{2,5}+e^{3,9},\\ &de^{14}=e^{1,5}+e^{4,9},\ \ de^{15}=e^{4,5}-e^{1,9},\ \ de^{16}=-e^{3,5}-e^{2,9}.\end{split}$
Then $\mathfrak{g}$ can be endowed with the following hypercomplex structure
$J_{1}e_{1}=e_{2},\ J_{1}e_{3}=e_{4},\ J_{1}e_{5}=e_{6},\ J_{1}e_{7}=e_{8},$
$J_{1}e_{9}=e_{10},\ J_{1}e_{11}=e_{12},\,J_{1}e_{13}=e_{14},\ J_{1}e_{15}=e_{16},$
$J_{2}e_{1}=e_{3},\ J_{2}e_{2}=-e_{4},\ J_{2}e_{5}=e_{7},\ J_{2}e_{6}=-e_{8},$
$J_{2}e_{9}=e_{11},\ J_{2}e_{10}=-e_{12},J_{2}e_{13}=e_{15},\ J_{2}e_{14}=-e_{16},$
$J_{3}e_{1}=e_{4},\ J_{3}e_{2}=e_{3},\ J_{3}e_{5}=e_{8},\ J_{3}e_{6}=e_{7},$
$J_{3}e_{9}=e_{12},\ J_{3}e_{10}=e_{11},\ J_{3}e_{13}=e_{16},\ J_{3}e_{14}=e_{15},$
satisfying $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, for each $\alpha=1,2,3$. Moreover, since $\mathfrak{g}$ is $3$-step, Proposition 7.2 implies that the Obata connection on the associated connected and simply connected nilpotent Lie group is not flat.
Theorem 6.1 and Proposition 7.2 highlight an unexpected rigidity in the holonomy algebra of the Obata connection. A related result is proved in [Ge], where the author shows that if the hypercomplex structure is abelian then the holonomy algebra of the Obata connection is abelian.
In the $3$-step nilpotent setting, as recalled above, all previously known examples were either Obata-flat or equipped with an abelian hypercomplex structure (the example constructed in this section is neither Obata-flat nor endowed with an abelian hypercomplex structure). Nevertheless, in every known case, the holonomy algebra of the Obata connection is abelian. This recurring behavior suggests that such rigidity might be a general feature of the nilpotent setting, motivating the following question:
Question 7.4.
Let $G$ be a nilpotent Lie group endowed with a left-invariant hypercomplex structure. Is it true that the holonomy algebra of the Obata connection is always abelian?